### v3

# Testing the Fairness-Accuracy Improvability of Algorithms

Abstract.
Many organizations use algorithms that have a disparate impact, i.e., the benefits or harms of the algorithm fall disproportionately on certain social groups. Addressing an algorithm’s disparate impact can be challenging, however, because it is often unclear whether it is possible to reduce this impact without sacrificing other objectives of the organization, such as accuracy or profit. Establishing the improvability of algorithms with respect to multiple criteria is of both conceptual and practical interest: in many settings, disparate impact that would otherwise be prohibited under US federal law is permissible if it is necessary to achieve a legitimate business interest. The question is how a policy-maker can formally substantiate, or refute, this “necessity” defense. In this paper, we provide an econometric framework for testing the hypothesis that it is possible to improve on the fairness of an algorithm without compromising on other pre-specified objectives. Our proposed test is simple to implement and can be applied under any exogenous constraint on the algorithm space. We establish the large-sample validity and consistency of our test, and illustrate its practical application by evaluating a healthcare algorithm originally considered by Obermeyer et al. (2019). In this application, we reject the null hypothesis that it is not possible to reduce the algorithm’s disparate impact without compromising the accuracy of its predictions.
We thank Denis Chetverikov, Joel Horowitz, Yiqi Liu, Francesca Molinari, Matthew Murphy, Azeem Shaikh and Eisho Takatsuji for helpful comments, and National Science Foundation Grants SES-1851629, SES-2149408 and SES-2149422 for financial support.
Department of Economics, Northwestern University
§Department of Economics, University of Chicago
1. Introduction
As algorithms are increasingly used to guide important predictions about people (e.g., which patients to treat, or which borrowers to grant a loan to), substantial concern has emerged about their potential disparate impact—namely, that the algorithm’s benefits or harms may be born unequally across different social groups. Disparate impact has been empirically documented in a range of applications (see for instance Angwin and Larson, 2016; Obermeyer et al., 2019; Arnold et al., 2021; Fuster et al., 2021), but the organizations that deploy these algorithms also value other objectives such as accuracy and profit. When an algorithm has a disparate impact, is it possible to reduce that disparity without compromising the organization’s other objectives? The answer to this question is legally relevant. In many settings, a policy with a disparate impact that would otherwise be prohibited under US federal law is permissible if it is necessary to achieve a legitimate business interest (see Section 1.1 for a discussion). As a result, it is essential for both external regulators and internal stakeholders to have tools that can determine when fairness is in conflict with other objectives, and when instead there exist alternative algorithms that do better on all of the specified criteria. In this paper, we provide a framework and results for establishing, or refuting, the existence of such algorithms. We also suggest practical approaches for finding these algorithms when they exist.
Our paper consists of three parts. In the first part, we formally define fairness-improvability, along with an accompanying notion of accuracy-improvability (where we use “accuracy” as an umbrella term for any objective not directly related to inequities across groups). In the second part, we provide a procedure for testing the fairness-improvability and accuracy-improvability of a status quo algorithm given data. Our proposed approach is simple to implement, flexible with respect to the definitions of fairness and accuracy, and can be applied under any exogenous constraint on the algorithm space. In the third part, we apply our procedure to show that a healthcare algorithm studied in Obermeyer et al. (2019) is both fairness- and accuracy-improvable.
To define fairness- and accuracy-improvability, we build on the framework of Liang et al. (2022). An algorithm is defined to be a mapping from a vector space of observed covariates (e.g., a medical profile) into a prediction (e.g., a medical diagnosis). There are two predefined groups. Algorithms are evaluated in terms of their accuracy for each group and their fairness across groups. Accuracy for each group is defined to be the group’s expected utility with respect to a pre-specified utility function, e.g., the algorithm’s correct classification rate or the fraction of patients in this group that get treated. Fairness is measured as the disparity in the expected utilities across groups, where the utility function used to evaluate fairness could be different from the one used to evaluate accuracy (but need not be). This very general formulation allows us to nest most objectives that have been proposed in the preceding literature on algorithmic fairness, as well as the broad range of business objectives that could be legally relevant (see Section 2.1 for examples).
Our analysis allows for the analyst to pre-specify a class of permissible algorithms. For any $\delta\in\mathbb{R}_{+}$, we define an algorithm to be $\delta$-fairness improvable within this class if some other algorithm in the class is at least as accurate as the first algorithm, and moreover strictly improves on its fairness by a factor of $\delta$. Similarly, we define an algorithm to be $\delta$-accuracy improvable in the class if some other algorithm in the class is at least as fair, and moreover strictly improves on accuracy by a factor of $\delta$. In both cases, improvability is defined with respect to a fixed population distribution.
In practice the analyst may not know this distribution. We suppose that the analyst does however have access to a sample of individuals, their covariates, and a “ground truth” outcome (i.e., a quantity that determines the optimal algorithm and is known ex-post). The sample can be used to estimate and make inferences about the fairness and accuracy of the status quo algorithm in the population. The goal of our test procedure is to infer whether there exists another algorithm that improves on both criteria simultaneously.
Our test procedure is as follows. We first split the data into two subsets: a training set and a test set. We then use the training set to identify a candidate algorithm that potentially improves on the status quo. Subject to certain regularity conditions, any method for identifying a candidate algorithm—which we call a selection rule—can be plugged into this step (see example selection rules from the literature in Section 1.2).
Once a candidate algorithm has been identified, we use the test set to evaluate whether the improvements of the candidate algorithm are statistically significant. To avoid computing standard errors case-by-case for each specification of accuracy and fairness, we propose using the bootstrap to compute the critical values of our test. While our use of the bootstrap is straightforward, its justification is complicated by two features of our problem. First, our test statistic is constructed using absolute values, and it is well-known that bootstrap consistency fails at points of non-differentiability (see e.g., Fang and Santos, 2019). We show that this is not an issue in the specific setting of our problem (see Section 3.2 for details). Second, the fact that we impose very few restrictions on the selection rule means that the distribution of our test statistics may vary with the sample size and are not guaranteed to “settle down” in the limit. To our knowledge, there does not exist a generally available bootstrap validity result that covers our exact setting. We thus prove a new bootstrap consistency result for our setting, building on a result for triangular arrays due to Mammen (1992): See Appendix B.2 for details.
To reduce the uncertainty introduced by sample-splitting, we recommend that the analyst repeats the above process several times and aggregates the results by reporting the median $p$-value across the set of tests. We show that an asymptotically valid test is obtained by rejecting the null hypothesis whenever the median $p$-value falls below half of the desired significance level. We further establish that this test is consistent under an additional assumption on the selection rule that we call improvement convergence. This condition says that whenever the null is not satisfied—i.e., the status quo algorithm is indeed $\delta$-fairness improvable (or $\delta$-accuracy improvable)—then the selection rule will asymptotically find an algorithm that attains this improvement over the status quo algorithm.
Our proposed approach overcomes a key practical challenge, which is that algorithms are often exogenously constrained for legal, logistical, or ethical reasons. The exact nature of these constraints may vary across settings, and include (among others) capacity constraints on how many individuals are admitted or treated, shape constraints such as monotonicity of the decision with respect to a given input (Wang and Gupta, 2020; Feelders, 2000; Liu et al., 2020), and statistical constraints such as requiring independence of decisions with a group identity variable (Mitchell et al., 2021; Barocas et al., 2023). Sample splitting and the bootstrap ensure that our test procedure is valid regardless of the specific class of algorithms, constraints, or utility functions chosen, subject to the regularity conditions given in Assumptions 1, 2 and 3.
Finally, to illustrate our proposed procedure, we revisit the data and setting of Obermeyer et al. (2019). The status quo algorithm is one used by a hospital to identify patients for automatic enrollment into a high-intensity care management program. Using our framework and test procedure, we reject the null hypothesis that it is not possible to simultaneously improve on the accuracy and the fairness of the algorithm. We further quantify the extent of possible improvement by testing the null that the hospital’s algorithm is not $\delta$-fairness improvable and $\delta$-accuracy improvable for different values of $\delta$. We find that large improvements in fairness are possible without compromising accuracy, while the reverse is not true.
1.1. Legal framework
The econometric framework we develop in this paper is broadly applicable whenever a policy-maker wants to examine whether it is possible to increase the fairness of a policy without sacrificing other objectives. However, we view our contribution as particularly relevant to regulators and other legal decision makers tasked with evaluating cases of alleged disparate impact.
To support this application, we review the legal process for making a disparate impact case under US federal law. [^1] We focus on restrictions of private individuals (as opposed to government agents) from certain policies. As a general rule, discrimination is legally permissible by private actors except when specifically prohibited by law. There is no singular discrimination law. However, there is a body of legislation, regulation, and court precedent that prohibits discrimination against certain social groups (such as those defined by race, gender, or religion) called “protected classes” in settings as varied as employment, healthcare, education, and lending. US federal law generally distinguishes between two types of discrimination: disparate treatment and disparate impact. A policy has a disparate treatment if it intentionally offers different services based on an individual’s membership in a protected class. It has a disparate impact if one protected class is benefited or harmed more than another by the policy. Our paper pertains to disparate impact, which we broadly review here. Details and specific references to the US Code, the Code of Federal Regulations, and the Department of Justice’s Guidelines are deferred to Appendix D.
A typical process for evaluating a disparate impact case has three parts. The first part is to determine whether the benefits or harms disproportionately affect members of a protected class. This is often called a “prima facie” case of discrimination. For this part of the process, there already exist well-established statistical frameworks (see for instance Hazelwood School District v. United States, 1977; Shoben, 1978; Groves v. Alabama State Bd. of Educ., 1991), so we do not focus on it in our paper.
If the prima facie case is not successful, then the process ends and concludes that there is no disparate impact. If it is successful, then the second part of the process is to determine whether the challenged policy is a business necessity, i.e., necessary to achieve some legitimate nondiscriminatory interest. If the business necessity defense is not successful, then the process ends and concludes that the disparate impact is not permissible. If it is successful, then the third and final part of the process is to determine whether there exists a valid alternative that is less discriminatory, but still serves the same interests as the challenged policy. The policy is not permissible if such an alternative exists. Otherwise, it is. Our econometric framework is designed to assess this final part.
The exact implementation of this three-part process depends on the setting, and in particular on the specific statute, regulation, or court precedent prohibiting discrimination that is allegedly violated. For example, in the context of employment discrimination (which is prohibited by Title VII of the Civil Rights Act of 1964), the Equal Employment Opportunity Commission and/or private individuals can bring a lawsuit to court to remedy a disparate impact violation. In this setting, the three-part test is adversarial; that is, there are two opposing parties, where one party is responsible for the evidence and arguments supporting the policy, and the other party is responsible for the evidence and arguments against the policy. In contrast, in the context of discrimination by a program or activity that receives federal funding (which is prohibited by Title VI of the Civil Rights Act of 1964), investigation and enforcement instead falls on the funding agency (see Alexander v. Sandoval, 2001). Though the funding agency’s determination of disparate impact is not typically viewed as adversarial, the Department of Justice still recommends using the three-part test outlined above. [^2] The Department of Justice is tasked with coordinating the implementation and enforcement of Title VI across federal agencies, by Executive Order 12250, 28 C.F.R. pt. 41, app. A. Their guidelines can be found in the Title VI Legal Manual available at https://www.justice.gov/crt/fcs/T6manual.
Twenty-six funding agencies currently enforce Title VI in a variety of contexts such as education, healthcare, and transportation. In Section 5, we apply our framework to study a potential disparate impact in access to hospital services. If the hospital receives federal funding from Medicare or Titles VI or XVI of the Public Health Service Act, this discrimination would be prohibited by Title VI of the Civil Rights Act of 1964. In this case, a likely enforcing agency would be the US Department of Health and Human Services (HHS). [^3] Specific referrals of individual complaints to HHS can be found at https://www.hhs.gov/civil-rights/for-individuals/index.html and https://www.cms.gov/about-cms/web-policies-important-links/accessibility-nondiscrimination-disabilities-notice. As hospitals and other healthcare providers increasingly use algorithms to allocate resources and determine patient care, we view our framework as one that could be used by HHS in order to evaluate allegations of discrimination under Title VI.
1.2. Related Literature
Our paper is related to several literatures in economics, statistics and computer science which study algorithmic fairness: see Chouldechova and Roth (2018), Cowgill and Tucker (2020), or Barocas et al. (2023) for recent overviews. Most of this literature is concerned with computational and algorithmic aspects of fair algorithm design. For example, many papers build an explicit fairness constraint into the optimization problem, and focus on the question of how to find (or approximate) the optimum in a computationally efficient manner (e.g., Dwork et al. (2012), Agarwal et al. (2018), Kusner et al. (2019), Dwork et al. (2018), Babii et al. (2020), and Viviano and Bradic (2023)). Coston et al. (2021) share our motivation of evaluating whether an algorithm with disparate impact can be justified by business necessity, and formulate an optimization problem in which disparate impact is minimized over a set of algorithms that satisfy an additional constraint.
We focus on the different but complementary problem of statistical inference, and develop an econometric framework to test whether there exists an algorithm that satisfies the same business objectives but has lower disparate impact. We abstract away from the question of how to efficiently search for a fairness-improving algorithm; any of the methods developed in this literature could be used within our proposed procedure as a means to identify candidate algorithms for fairness-improvement.
Our sample-splitting procedure is closely related to the procedure proposed in Blattner and Spiess (2023), which studies potential fairness and interpretability improvements in lending algorithms. Although the mechanics of our two procedures are different, we both use training data to identify a “less discriminatory algorithm” and (disjoint) test data to evaluate it. Relative to this work, our paper contributes formal statistical guarantees.
Another closely related paper is the work of Liu and Molinari (2024), which directly estimates the fairness-accuracy frontier developed by Liang et al. (2022). A full characterization of the frontier allows the analyst to answer a broad range of questions related to algorithmic fairness. In particular, once the fairness-accuracy frontier has been characterized, testing for the fairness-improvability of a status-quo algorithm amounts to testing whether or not the algorithm belongs to the frontier. While our approach focuses on the narrower question of testing for fairness-improvability (rather than characterizing the full frontier), it is more flexible in two ways—(1) it allows for definitions of accuracy and fairness with respect to different utility functions, and (2) can automatically accommodate (almost) any exogenous constraint on the algorithm space. [^4] The characterization of the frontier in Liu and Molinari (2024) is developed for the class of all possible algorithms, and (to the best of our knowledge) holds only under specific restrictions on the algorithm space, which rule out global constraints such as capacity constraints and monotonicity constraints. Both of these features are relevant for evaluating the kinds of algorithms that are used in practice.
2. Model
2.1. Setup
Consider a population of individuals, where each individual is described by a covariate vector $X$ taking values in the set $\mathcal{X}\subseteq\mathbb{R}^{d}$, a type $Y$ taking values in $\mathcal{Y}\subseteq\mathbb{R}$, and a group identity $G$ taking values in $\mathcal{G}=\{r,b\}$. The covariates in $X$ are the information that the algorithm can access to make a decision, the type $Y$ is a payoff-relevant unknown that the algorithm cannot directly access, and the group identity $G$ is a special covariate (which the algorithm may or may not access [^5] The algorithm is given only $X$ as input, but we allow for the possibility that $G$ is a covariate in $X$, or is perfectly correlated with covariates in $X$.) that is used to evaluate the fairness of the algorithm. We use $P$ to denote the joint distribution of $(X,Y,G)$ across individuals. An algorithm $a:\mathcal{X}\rightarrow\mathcal{D}$ maps covariate vectors into $\mathcal{D}$, the set of possible decisions.
We evaluate algorithms from the perspectives of accuracy and fairness, which are defined with respect to an accuracy utility function $u_{A}:\mathcal{X}\times\mathcal{Y}\times\mathcal{D}\rightarrow\mathbb{R}$, and a (possibly identical) fairness utility function $u_{F}:\mathcal{X}\times\mathcal{Y}\times\mathcal{D}\rightarrow\mathbb{R}$. We will consider accuracy and fairness criteria that can be formulated as
(2.1) $U_{A}^{g}(a)=\frac{E_{P}[u_{A}(X,Y,a(X))\mid G=g]}{E_{P}[w_{A}(X,Y,a(X))\mid G=g]}\quad\mbox{and}\quad U_{F}^{g}(a)=\frac{E_{P}[u_{F}(X,Y,a(X))\mid G=g]}{E_{P}[w_{F}(X,Y,a(X))\mid G=g]}$
where $w_{A}$ and $w_{F}$ are normalization functions (see Examples 1-4 for possible specifications). Thus, $U_{A}^{g}$ and $U_{F}^{g}$ are normalized expected utilities for each group according to the respective utility functions. When the utility functions and normalization functions are common across accuracy and fairness, we drop the subscripts and simply write $U^{g}(a)$.
We say that one algorithm is more accurate than another if the value of $U_{A}^{g}$ is larger for both groups, and more fair if the value of $U^{g}_{F}$ differs less across groups.
Definition 2.1.
Algorithm $a_{1}$ is more accurate than algorithm $a_{0}$ if
$U_{A}^{r}(a_{1})>U_{A}^{r}(a_{0})\quad\mbox{and}\quad U_{A}^{b}(a_{1})>U_{A}^{b}(a_{0})$
and more fair than algorithm $a_{0}$ if
$|U_{F}^{r}(a_{1})-U_{F}^{b}(a_{1})|<|U_{F}^{r}(a_{0})-U_{F}^{b}(a_{0})|.$
We list below example utility specifications from the prior literature. [^6] Examples 1-3 are popular in the algorithmic fairness literature (see e.g. Mitchell et al. (2021)), while Example 4 is a canonical economic example. To simplify exposition, we drop the fairness and accuracy subscripts, understanding that the designer could use different utility specifications for the two criteria.
Example 1 (Classification Rate).
The decision is a prediction of $Y$, and both are binary. Let $u(x,y,d)=\mathbbm{1}(y=d)$ be an indicator variable for whether the prediction is correct, and let the normalization function be $w(x,y,d)=1$. Then
$U^{g}(a)=P(Y=a(X)\mid G=g)$
is the correct classification rate of algorithm $a$ within group $g$, while $|U^{r}(a)-U^{b}(a)|$ is the disparity in classification rates across groups.
Example 2 (Calibration).
The type $Y$ takes values in $\mathbb{R}$ and the decision is binary. Define $u(x,y,d)=y\mathbbm{1}(d=1)$ and $w(x,y,d)=\mathbbm{1}(d=1)$. Then
$U^{g}(a)=E[Y\mid a(X)=1,G=g]$
is the expected type for individuals receiving decision $a(X)=1$ in group $g$, while $|U^{r}(a)-U^{b}(a)|$ evaluates how different these conditional expectations are across groups.
Example 3 (False Positive Rate).
Both $Y$ and the decision are binary. Define $u(x,y,d)=\mathbbm{1}(y=0,d=1)$ and $w(x,y,d)=\mathbbm{1}(y=0)$. Then
$U^{g}(a)=P(a(X)=1\mid Y=0,G=g)$
is the probability of choosing the (wrong) decision $a(X)=1$ for individuals with true type $Y=0$ in group $g$, while $|U^{r}(a)-U^{b}(a)|$ compares the probability of this error across groups.
Example 4 (Profit).
There is a good that is offered to an individual at price $d$. The individual’s unknown willingness-to-pay is $Y$. The utility function is
$u(x,y,d)=\left\{\begin{array}[]{cc}d&\mbox{ if }y\geq d\\ 0&\mbox{ otherwise}\end{array}\right.$
and the normalization function is $w(x,y,d)=1$. That is, the firm receives the price if the individual’s willingness to pay exceeds that price, and otherwise the firm receives nothing. Then $U^{g}(a)$ is the average profit the firm receives from group $g$ consumers, and $|U^{r}(a)-U^{b}(a)|$ is the difference in average profit between consumers in the two groups.
In some applications, the analyst may wish to consider an accuracy measure that is unconditional across groups (e.g., the expected utility in the entire population), or to require simultaneous fairness improvements according to several different utility functions. [^7] For example, the analyst may require a fairness improvement not only with respect to the utility function given in Example 2, but also with respect to the mirrored utility function that conditions on $D=0$. Indeed, the property popularly known as calibration requires improvements according to both of these measures (Chouldechova, 2017; Kleinberg et al., 2017). Our procedure and results can be straightforwardly extended in either of these cases, but we will present our results for the three criteria $U_{A}^{r}$, $U_{A}^{b}$, and $|U_{F}^{r}-U_{F}^{b}|$ proposed here.
For a fixed class of algorithms $\mathcal{A}$, the feasible accuracy and unfairness levels are those $(U_{A}^{r},U_{A}^{b},|U_{F}^{r}-U_{F}^{b}|)$ triples that can be achieved by some algorithm in $\mathcal{A}$. Our proposed testing procedure will be agnostic about the specific choice of $\mathcal{A}$, and in particular we will allow for $\mathcal{A}$ to be exogenously restricted in some way. We view this flexibility as important since (as discussed in the introduction) in practice there are often legal, logistical, or ethical constraints on which algorithms can be implemented.
2.2. Accuracy/Fairness Improvability
There is a status quo algorithm $a_{0}:\mathcal{X}\rightarrow\mathcal{D}$ representing an algorithm that is already in use, or which has been proposed for use. The status quo algorithm $a_{0}$ has some level of unfairness, $|U_{F}^{r}(a_{0})-U_{F}^{b}(a_{0})|$, which we assume throughout the paper is strictly positive.
Assumption 1.
The status quo algorithm is not perfectly fair, i.e., $|U_{F}^{r}(a_{0})-U_{F}^{b}(a_{0})|>0$.
This assumption reflects the basic motivation of our paper: If the status quo algorithm has no disparate impact (corresponding to $|U_{F}^{r}(a_{0})-U_{F}^{b}(a_{0})|=0$), then there is no reason to look for fairness-improving alternatives. Supposing that Assumption 1 holds, we ask whether it is possible to improve upon the disparate impact of the status quo algorithm without compromising on its accuracy. Formally, we extend Definition 2.1 in the following way to quantify the extent of fairness- or accuracy-improvement.
Definition 2.2 (Accuracy-Fairness Improvement).
Fix any $\Delta_{r},\Delta_{b},\Delta_{f}\in\mathbb{R}_{+}$. The algorithm $a_{1}$ constitutes a $(\Delta_{r},\Delta_{b},\Delta_{f})$-improvement on the algorithm $a_{0}$ if
$\frac{U_{A}^{r}(a_{1})}{U_{A}^{r}(a_{0})}>1+\Delta_{r},\quad\frac{U_{A}^{b}(a_{1})}{U_{A}^{b}(a_{0})}>1+\Delta_{b},\quad\mbox{and}\quad\frac{|U_{F}^{r}(a_{1})-U_{F}^{b}(a_{1})|}{|U_{F}^{r}(a_{0})-U_{F}^{b}(a_{0})|}<1-\Delta_{f}~{}.$
That is, algorithm $a_{1}$ constitutes a $(\Delta_{r},\Delta_{b},\Delta_{f})$-improvement on algorithm $a_{0}$ if it increases accuracy for group $r$ by a factor of $\Delta_{r}$, increases accuracy for group $b$ by a factor of $\Delta_{b}$, and reduces disparate impact by a factor of $\Delta_{f}$. Analogous results to those presented in Sections 3 and 4 hold under a modification of Definition 2.2, where we replace the ratios in this definition with differences. [^8] There are advantages and disadvantages to each of these comparisons: Comparing expected utilities by their ratios is invariant to scaling of utilities, but sensitive to constant translations, while comparing expected utilities by their difference is invariant to constant translations and sensitive to scaling. We expect that in applications of our framework, the functions $U$ will not be an abstract measure of utility, but rather an explicit measure of something concrete (such as in our examples in Section 2.1 and 5), rendering invariances to transformations of these forms less important.
Below, we give names to special cases that we want to emphasize.
Definition 2.3 (FA-dominance).
Fix a class of algorithms $\mathcal{A}$. The algorithm $a_{0}$ is FA-dominated within class $\mathcal{A}$ if there exists an algorithm $a_{1}\in\mathcal{A}$ that $(0,0,0)$-improves on $a_{0}$, i.e.,
$\frac{U_{A}^{r}(a_{1})}{U_{A}^{r}(a_{0})}>1,\quad\frac{U_{A}^{b}(a_{1})}{U_{A}^{b}(a_{0})}>1,\quad\mbox{and}\quad\frac{|U_{F}^{r}(a_{1})-U_{F}^{b}(a_{1})|}{|U_{F}^{r}(a_{0})-U_{F}^{b}(a_{0})|}<1~{}.$
If an algorithm is FA-dominated, then it is possible to simultaneously strictly improve upon its accuracy (for both groups) as well as its fairness. We view this definition as directly related to the business-necessity defense in a disparate impact case (as described in Section 1.1). When $U_{A}^{g}$ captures a business-relevant objective, then FA-dominance corresponds to the existence of an alternative algorithm that can achieve the same business objectives with less disparate impact.
We can further quantify the extent to which it is possible to improve either fairness or accuracy, without compromising on the other objective.
Definition 2.4 ($\delta$-Fairness Improvability).
Fix a class of algorithms $\mathcal{A}$ and any $\delta\in\mathbb{R}_{+}$. The algorithm $a_{0}$ is $\delta$-fairness improvable within class $\mathcal{A}$ if there exists an algorithm $a_{1}\in\mathcal{A}$ that $(0,0,\delta)$-improves on $a_{0}$, i.e.,
$\frac{U_{A}^{r}(a_{1})}{U_{A}^{r}(a_{0})}>1,\quad\frac{U_{A}^{b}(a_{1})}{U_{A}^{b}(a_{0})}>1,\quad\mbox{and}\quad\frac{|U_{F}^{r}(a_{1})-U_{F}^{b}(a_{1})|}{|U_{F}^{r}(a_{0})-U_{F}^{b}(a_{0})|}<1-\delta~{}.$
If an algorithm is $\delta$-fairness improvable, then it is possible for the firm to strictly improve on fairness by at least a factor of $\delta$ without compromising on accuracy.
Definition 2.5 ($\delta$-Accuracy Improvability).
Fix a class of algorithms $\mathcal{A}$ and any $\delta\in\mathbb{R}_{+}$.
  • (a)
    The algorithm $a_{0}$ is $\delta$-accuracy improvable for group $g\in\{r,b\}$ within class $\mathcal{A}$ if there exists an algorithm $a_{1}\in\mathcal{A}$ that $(\delta,0,0)$-improves on $a_{0}$, i.e.,
    $\frac{U_{A}^{g}(a_{1})}{U_{A}^{g}(a_{0})}>1+\delta,\quad\frac{U_{A}^{g^{\prime}}(a_{1})}{U_{A}^{g^{\prime}}(a_{0})}>1,\quad\mbox{and}\quad\frac{|U_{F}^{g}(a_{1})-U_{F}^{g^{\prime}}(a_{1})|}{|U_{F}^{g}(a_{0})-U_{F}^{g^{\prime}}(a_{0})|}<1~{}.$
    where $g^{\prime}\neq g$.
  • (b)
    The algorithm $a_{0}$ is $\delta$-jointly accuracy improvable within class $\mathcal{A}$ if there exists an algorithm $a_{1}\in\mathcal{A}$ that $(\delta,\delta,0)$-improves on $a_{0}$, i.e.,
    $\frac{U_{A}^{r}(a_{1})}{U_{A}^{r}(a_{0})}>1+\delta,\quad\frac{U_{A}^{b}(a_{1})}{U_{A}^{b}(a_{0})}>1+\delta,\quad\mbox{and}\quad\frac{|U_{F}^{r}(a_{1})-U_{F}^{b}(a_{1})|}{|U_{F}^{r}(a_{0})-U_{F}^{b}(a_{0})|}<1~{}.$
That is, an algorithm is $\delta$-accuracy improvable for group $g$ if it is possible to improve accuracy for group $g$ by at least a factor of $\delta$ without compromising on accuracy for the other group or on fairness. The algorithm is $\delta$-jointly accuracy improvable if it is possible to improve accuracy for both groups by at least a factor of $\delta$ without compromising on fairness. These definitions coincide and reduce to FA-dominance in the special case of $\delta=0$.
3. Proposed Approach
We consider an analyst who does not know the joint distribution of $(X,Y,G)$, but instead observes a sample $\{(X_{i},Y_{i},G_{i})\}_{i=1}^{n}$ consisting of $n$ independent and identically distributed observations from this distribution. The analyst’s objective is to test fairness-improvability and accuracy-improvability of a status quo algorithm $a_{0}$ given such a sample. Formally, the analyst specifies a set of algorithms $\mathcal{A}$ and tests the null hypothesis
(3.1) $H_{0}:\text{Algorithm $a_{0}$ is not $\delta$-fairness (or $\delta$-accuracy) improvable within class $\mathcal{A}$},$
against the alternative that such an improvement exists. Section 3.1 outlines the approach, and the subsequent Section 4 establishes the validity and consistency of our procedure under stated assumptions.
3.1. Description of Procedure
The analyst first chooses a selection rule that maps samples into a choice of algorithm from $\mathcal{A}$, i.e., a mapping
$\rho:\mathcal{S}\rightarrow\mathcal{A}$
where $\mathcal{S}=\bigcup_{m\geq 1}\mathcal{S}_{m}\equiv\bigcup_{m\geq 1}(\mathcal{X}\times\mathcal{Y}\times\mathcal{G})^{m}$ is the set of all finite samples of observations. Our results apply for any such rule subject to certain regularity conditions which we introduce in Section 3.2.
In each round $k=1,\dots,K$ of the procedure (where $K$ is chosen by the user), the analyst repeats the following steps:
[1] Split the sample into train and test. The training sample $S_{train}^{n}$ includes $m_{n}=\lfloor\beta\cdot n\rfloor$ observations selected uniformly at random from $\{(X_{i},Y_{i},G_{i})\}_{i=1}^{n}$, where $\beta\in[0,1]$ is a user-chosen fraction (we use $\beta=1/2$ in our application). The test sample $S_{test}^{n}$ consists of the remaining $\ell_{n}=n-m_{n}$ observations.
[2] Find a candidate algorithm using the training sample. Apply the selection rule $\rho$ to the training sample $S_{train}^{n}$. To simplify notation, we use $\hat{a}^{\rho}_{1n}$ to denote the (realized) algorithm $\rho(S_{train}^{n})$, which serves as a candidate algorithm in the subsequent step.
[3] Test whether $\hat{a}^{\rho}_{1n}$ constitutes a $(\Delta_{r},\Delta_{b},\Delta_{f})$-improvement relative to $a_{0}$. Given the status quo algorithm $a_{0}$ and the candidate algorithm $\hat{a}^{\rho}_{1n}$ obtained in the previous step, test the null hypothesis
(3.2) $H_{0}:\text{$\hat{a}^{\rho}_{1n}$ does not constitute a $(\Delta_{r},\Delta_{b},\Delta_{f})$-improvement on $a_{0}$}$
using the test outlined in Section 3.2 and the sample $S_{test}^{n}$. We set $(\Delta_{r},\Delta_{b},\Delta_{f})=(0,0,0)$ in a test of FA-dominance, $(\Delta_{r},\Delta_{b},\Delta_{f})=(0,0,\delta)$ in a test of $\delta$-fairness improvability, $(\Delta_{r},\Delta_{b},\Delta_{f})=(\delta,0,0)$ in a test of $\delta$-accuracy improvability for group $r$, and $(\Delta_{r},\Delta_{b},\Delta_{f})=(\delta,\delta,0)$ in a test of $\delta$-joint accuracy improvability. Let $p_{k}$ denote the $p$-value associated with this test (details below in Section 3.2).
[4] Repeat steps 1-3 $K$ times, yielding a vector of $p$-values $(p_{1},\dots,p_{K})$.
[5] Define $p_{\mathrm{med}}$ to be the median of the values $(p_{1},\dots,p_{K})$ and reject $H_{0}$ if and only if $p_{\mathrm{med}}<\alpha/2$. Corollary 4.1 guarantees that under certain regularity conditions, this test is of level $\alpha$. This proposed procedure is summarized in Figure 1.
Figure 1. A summary of our procedure for testing FA improvability.
3.2. Testing whether $\hat{a}^{\rho}_{1n}$ constitutes a $(\Delta_{r},\Delta_{b},\Delta_{f})$-improvement on $a_{0}$.
This section provides supporting details and results for step [3] of our procedure outlined in Section 3.1. Since $\hat{a}^{\rho}_{1n}$ was selected using the training sample $S_{train}$, it suffices to describe a test of $(\Delta_{r},\Delta_{b},\Delta_{f})$-improvement for a fixed algorithm $a_{1}$ relative to $a_{0}$, using the test sample $S_{test}^{n}=\{(X_{i},Y_{i},G_{i})\}_{i=1}^{\ell_{n}}$ where $\ell_{n}=n-m_{n}$. To that end, we test the null hypothesis
(3.3) $H_{0}:A_{1r}\leq A_{0r}(1+\Delta_{r})\text{ OR }A_{1b}\leq A_{0b}(1+\Delta_{b})\text{ OR }|F_{1r}-F_{1b}|\geq|F_{0r}-F_{0b}|(1-\Delta_{f})~{},$
against the alternative
$H_{1}:A_{1r}>A_{0r}(1+\Delta_{r})\text{ AND }A_{1b}>A_{0b}(1+\Delta_{b})\text{AND }|F_{1r}-F_{1b}|<|F_{0r}-F_{0b}|(1-\Delta_{f})~{},$
for an arbitrary vector $\Delta\equiv(\Delta_{r},\Delta_{b},\Delta_{f})\in\mathbb{R}^{3}_{+}$, where
$(A_{0r},A_{0b},F_{0r},F_{0b})\equiv(U_{A}^{r}(a_{0}),U_{A}^{b}(a_{0}),U_{F}^{r}(a_{0}),U_{F}^{b}(a_{0}))$
denotes the accuracy and fairness levels under the status quo algorithm $a_{0}$ and
$(A_{1r},A_{1b},F_{1r},F_{1b})\equiv(U_{A}^{r}(a_{1}),U_{A}^{b}(a_{1}),U_{F}^{r}(a_{1}),U_{F}^{b}(a_{1}))$
denotes the accuracy and fairness levels under the candidate algorithm $a_{1}$.
Since $H_{0}$ is formulated as a union of three conditions, a test of $H_{0}$ can be constructed by first constructing tests for each of the conditions individually, and then combining the individual tests using the intersection-union method. [^9] See Casella and Berger (2021) for a textbook reference on the intersection-union method. Proposition B.1 in the appendix extends Theorems 8.3.23-24 in Casella and Berger (2021) to an asymptotic setting. Specifically, we will construct tests for each of the individual conditions that make up $H_{0}$, and reject $H_{0}$ if all three tests reject at once.
First, we specify test statistics for each of the individual tests that make up $H_{0}$. To simplify the derivations, we introduce a slight change in notation. Let $Z_{i}=(X_{i},Y_{i},G_{i})$ and define
$u_{tg}^{A}(Z_{i},\theta)$ $=\frac{1}{\theta^{A}_{tg}}\left(u_{A}(X_{i},Y_{i},a_{t}(X_{i}))\cdot\mathbbm{1}(G_{i}=g)\right)$
$u_{tg}^{F}(Z_{i},\theta)$ $=\frac{1}{\theta^{F}_{tg}}\left(u_{F}(X_{i},Y_{i},a_{t}(X_{i}))\cdot\mathbbm{1}(G_{i}=g)\right)$
where
$\theta^{A}_{tg}$ $=E\left[w_{A}(X_{i},Y_{i},a_{t}(X_{i}))\cdot\mathbbm{1}(G_{i}=g)\right]$
$\theta_{tg}^{F}$ $=E\left[w_{F}(X_{i},Y_{i},a_{t}(X_{i}))\cdot\mathbbm{1}(G_{i}=g)\right]$
The vector $\theta\in\mathbb{R}^{J}$ collects all of the unknown nuisance parameters $(\theta^{A}_{tg},\theta^{F}_{tg})$ which appear as normalizations in the utility specifications. This re-writing allows for the accuracy and fairness criteria $A_{tg}$ and $F_{tg}$ to be expressed as unconditional expectations of the functions $u_{tg}$, rather than as ratios of the conditional expectations of $u_{A}$, $u_{F}$, $w_{A}$, $w_{F}$ (as defined in (2.1)). That is, $A_{tg}=E[u^{A}_{tg}(Z_{i},\theta)]$ and $F_{tg}=E[u^{F}_{tg}(Z_{i},\theta)]$.
In the results that follow, we will in fact allow $u^{A}_{tg}$ and $u^{F}_{tg}$ to depend on the nuisance parameter $\theta$ more generally than what is considered above (see specifically Assumption 2). Since $\theta$ is unknown, it will need to be estimated. We assume that $\theta$ can be written as $\theta=E[h_{i}]$ where $h_{i}=h(Y_{i},G_{i},X_{i},a_{1}(X_{i}),a_{0}(X_{i}))$ for some known function $h:\mathcal{Y}\times G\times\mathcal{X}\times\mathcal{D}^{2}\to\mathbb{R}^{J}$. This characterizes the four examples from Section 2. For instance, consider the following example based on the calibration utility in Example 2.
Example 5.
Suppose we wish to test the null hypothesis $H_{0}$ in a setting where $u_{A}=u_{F}=u$, and $u$ is given by the calibration utility defined in Example 2. Then $u^{A}_{tg}=u^{F}_{tg}=u_{tg}$ is given by
$u_{tg}(Z_{i},\theta)=\frac{1}{\theta_{tg}}Y_{i}\mathbbm{1}\{a_{t}(X_{i})=1\}\mathbbm{1}\{G_{i}=g\}$
where $\theta=(\theta_{0r},\theta_{1r},\theta_{0b},\theta_{1b})=E[h(G_{i},a_{1}(X_{i}),a_{0}(X_{i})]$ and
$h(G_{i},a_{1}(X_{i}),a_{0}(X_{i}))=\left(\begin{array}[]{c}\mathbbm{1}(G_{i}=r,a_{0}(X_{i})=1)\\ \mathbbm{1}(G_{i}=r,a_{1}(X_{i})=1)\\ \mathbbm{1}(G_{i}=b,a_{0}(X_{i})=1)\\ \mathbbm{1}(G_{i}=b,a_{1}(X_{i})=1)\end{array}\right)~{}.$
We estimate $A_{tg}$, $F_{tg}$, and $\theta$ using their empirical analogs $\hat{A}_{tg}=\frac{1}{\ell_{n}}\sum_{i}u^{A}_{tg}(Z_{i},\hat{\theta})$, $\hat{F}_{tg}=\frac{1}{\ell_{n}}\sum_{i}u^{F}_{tg}(Z_{i},\hat{\theta})$ and $\hat{\theta}=\frac{1}{\ell_{n}}\sum_{i}h_{i}$. Our test statistics for the individual tests are then defined as follows. When testing whether $a_{1}$ is less accurate than $a_{0}$ for groups $r$ and $b$, let
$\hat{T}_{r,n}\equiv\sqrt{\ell_{n}}(\hat{A}_{1r}-(1+\Delta_{r})\hat{A}_{0r})~{},$
$\hat{T}_{b,n}\equiv\sqrt{\ell_{n}}(\hat{A}_{1b}-(1+\Delta_{b})\hat{A}_{0b})~{}.$
When testing whether $a_{1}$ is less fair than $a_{0}$, define
$\hat{T}_{f,n}\equiv\sqrt{\ell_{n}}\left(\big{|}\hat{F}_{1r}-\hat{F}_{1b}\big{|}-(1-\Delta_{f})\big{|}\hat{F}_{0r}-\hat{F}_{0b}\big{|}\right)$
To avoid having to compute appropriate standard errors on a case-by-case basis for each utility function, we instead propose using the nonparametric bootstrap to generate the critical values of our test. Specifically, if we fix $\alpha$ at some nominal level, then for the accuracy parts of the test, the rejection rule is given by $\phi_{n}^{(g)}=\mathbbm{1}\{\hat{T}_{g,n}>c^{*}_{g,1-\alpha}\}$ for $g\in\{r,b\}$, where $c^{*}_{g,1-\alpha}=\Psi_{g,n}^{-1}(1-\alpha)$ is the $(1-\alpha)$-quantile of the bootstrap distribution $\Psi_{g,n}$ defined in the appendix. For the fairness part of the test, the rejection rule is given by $\phi_{n}^{(f)}=\mathbbm{1}\{\hat{T}_{f,n}<c^{*}_{f,\alpha}\}$ where $c^{*}_{f,\alpha}=\Psi_{f,n}^{-1}(\alpha)$ is the $\alpha$-quantile of the bootstrap distribution $\Psi_{f,n}$ defined in Appendix A. Correspondingly, the $p$-value associated with the test $\phi_{n}^{(g)}$ is given by $1-\Psi_{g,n}(\hat{T}_{g,n})$ and the $p$-value associated with the test $\phi_{n}^{(f)}$ is given by $\Psi_{f,n}(\hat{T}_{f,n})$. Our test of the null hypothesis (3.3) is finally given by
$\phi_{n}(a_{0},a_{1})=\phi_{n}^{(r)}\cdot\phi_{n}^{(b)}\cdot\phi_{n}^{(f)},$
so we reject if and only if all three component tests reject. As a result, the $p$-value associated with the test $\phi_{n}$ is given by the maximum of the three $p$-values associated with each of the component tests.
The validity of our bootstrap procedure hinges crucially on the structure of our specific problem. Indeed, given the form of $\hat{T}_{f,n}$, it is well known that a standard nonparametric bootstrap fails to be consistent at points of non-differentiability (see for instance Fang and Santos, 2019). This turns out not to be a problem for establishing the validity of our test for the following reason: recall that the null is a union of three component null hypotheses. If the null holds because $A_{1g}\leq A_{0g}(1+\Delta_{g})$ for some $g\in\{r,b\}$, then (as we show in the proof of Theorem 4.1) it is not necessary for the fairness test $\phi_{n}^{(f)}$ to control size, since one of the other two tests will. Thus, the only region of the null space where the size of $\phi_{n}^{(f)}$ is relevant is in the region where
$|F_{1r}-F_{1b}|\geq|F_{0r}-F_{0b}|(1+\Delta_{f})~{},$
but Assumption 1 implies that the RHS is strictly positive, so also $|F_{1r}-F_{1b}|>0$. As a result, we will be bounded away from points of non-differentiability of $\hat{T}_{f,n}$ in this region of the null space. [^10] Under the alternative, we prove consistency of the bootstrap under a mirrored assumption that the selection rule selects candidate algorithms which are not arbitrarily fair; we return to this point in the discussion following Theorem 4.2.
4. Main Results
We now show that under certain regularity conditions the test proposed in Section 3 is asymptotically valid when the candidate algorithm $a_{1}$ is selected using the data in $S^{n}_{train}$. We will also show that the test is consistent under an additional assumption on the selection rule $\rho(\cdot)$ which we call improvement convergence.
First consider $K=1$, i.e., there is only one round of sample-splitting. Let $(S_{train}^{n},S_{test}^{n})$ denote the realized split of training and testing samples of the data $\{(X_{i},Y_{i},G_{i})\}_{i=1}^{n}$. Denote the candidate algorithm selected using rule $\rho(\cdot)$ on the training sample $S_{train}^{n}$ as
$\hat{a}^{\rho}_{1n}\equiv\rho(S_{train}^{n}).$
Consider using the test statistic $\phi_{n}(a_{0},\hat{a}^{\rho}_{1n})$ (as defined in Section 3.2) to test the null hypothesis (3.3) on the test sample $S_{test}^{n}$. We first establish the asymptotic validity of the test under the following regularity conditions restricting the distribution $P$ and class of algorithms $\mathcal{A}$:
Assumption 2.
The nuisance parameters $\theta$ belong to an open convex subset $\Theta\subset\mathbb{R}^{J}$. The functions $u_{tg}^{A}(z,\cdot)$ and $u_{tg}^{F}(z,\cdot)$ are twice continuously differentiable on $\Theta$ for every $z\in\mathcal{Z}\equiv\mathcal{X}\times\mathcal{Y}\times\mathcal{G}$, $t\in\{0,1\}$, and $g\in\{r,b\}$. Furthermore, $h(\cdot)$ is bounded, and $u_{tg}^{A}$, $u_{tg}^{F}$, and their first and second-order partial derivatives (with respect to $\theta$) are all uniformly bounded over $(z,\theta)\in\mathcal{Z}\times\Theta$.
Assumption 3.
Let $\mathcal{M}$ denote the set of covariance matrices of the nuisance parameters and utilities (as defined in (B.3) in the appendix) for every candidate algorithm $a_{1}\in\mathcal{A}$. Then we assume
$\inf_{\Sigma\in\mathcal{M}}\lambda_{\rm min}(\Sigma)\geq\nu~{},$
for some $\nu>0$, where $\lambda_{\rm min}(\Sigma)$ denotes the smallest eigenvalue of $\Sigma$.
Assumption 2 allows us to appropriately linearize the utility functions as a function of $\theta$, uniformly over the parameter space $\Theta$. The uniformity of the linearization is important since the true value of the nuisance parameter can fluctuate as a function of sample size. The boundedness conditions in Assumption 2 impose implicit constraints on the parameter space $\Theta$ and the support of the data. For instance, revisiting Example 5, we see that the support of $Y_{i}$ should be bounded, and each component of $\theta$ should be bounded away from zero. Assumption 3 ensures that the limiting variances of our test statistics are non-degenerate uniformly in the space of candidate algorithms $\mathcal{A}$. Once again, this uniform non-degeneracy is important because the candidate algorithm is allowed to vary arbitrarily as a function of sample size. Low-level conditions which guarantee Assumption 3 are difficult to articulate at this level of generality, but Assumption 3 intuitively requires that the variances of the utilities be bounded away from zero uniformly in the space of candidate algorithms, and that the correlation in the utilities between the status-quo and candidate algorithms be bounded away from one. We expect these conditions to be satisfied for most selection rules used in practice. [^11] Note that, more precisely, we in fact only require Assumption 3 to hold for the set of algorithms $\mathcal{A}^{\prime}\subseteq\mathcal{A}$ which can be realized by the selection rule $\rho(\cdot)$.
Theorem 4.1 establishes that the test $\phi_{n}$ is asymptotically of level $\alpha$ regardless of the choice of algorithms $\hat{a}^{\rho}_{1n}$, as long as the nuisance parameters and utilities are well-behaved over the class $\mathcal{A}$ as described in Assumptions 1, 2 and 3.
Theorem 4.1.
Suppose $P$ and $\mathcal{A}$ satisfy Assumptions 1, 2 and 3, and suppose the null hypothesis given in (3.1) holds. Then
$\limsup_{n\rightarrow\infty}E_{P}[\phi_{n}(a_{0},\hat{a}^{\rho}_{1n})]\leq\alpha~{}.$
In practice, a researcher may prefer to combine the test results over multiple splits of the data to reduce the unpredictability in the procedure induced by sample splitting. Formally, consider $K>1$ and let $(S_{train}^{n,k},S_{test}^{n,k})$ denote the realized split of training and testing samples in the $k$-th round of sample splitting. Further let
$\hat{a}^{\rho}_{1nk}\equiv\rho(S_{train}^{n,k})$
denote the candidate algorithm selected using rule $\rho$ on the training sample $S_{train}^{n,k}$. Define the test statistic $\psi_{n}=\mathbbm{1}\left\{\frac{1}{K}\sum_{k=1}^{K}\phi_{n,k}(a_{0},\hat{a}^{\rho}_{1nk})\geq 1/2\right\}$ to reject if the median test statistic (across the rounds of sample splitting) rejects. We establish the following corollary of Theorem 4.1:
Corollary 4.1.
Under the assumptions of Theorem 4.1,
$\limsup_{n\rightarrow\infty}E_{P}[\psi_{n}]\leq 2\alpha~{}.$
Note that, as a consequence of Corollary 4.1, to ensure that the test $\psi_{n}$ is asymptotically of level $\alpha$ we require that the component tests $\phi_{n,k}$ for $k=1,\ldots,K$ are of level $\alpha/2$. Equivalently, we reject the null if the median $p$-value (across the rounds of sample splitting) satisfies $p_{\mathrm{med}}<\alpha/2$. We would in general expect this conservativeness to result in a loss of power relative to a test performed using a single split, but the local power analysis conducted in DiCiccio et al. (2020) suggests that this may not always be the case for alternatives not too close to the null; in particular, their Theorem 3.1 demonstrates that this is the case in a one-sided test of the mean, which is similar to the component nulls in (3.3).
Finally, we establish the consistency of the test $\phi_{n}(a_{0},\hat{a}^{\rho}_{1n})$ under high-level conditions on the behavior of the selection rule $\rho$.
Definition 4.1.
[…]
In words, improvement convergence says that whenever there exists a fairness (or accuracy) improvement on the algorithm $a_{0}$ in $\mathcal{A}$, the selection rule $\rho$ is guaranteed to find it asymptotically. Which of $\delta$-fairness or $\delta$-accuracy improvement is used in the definition should correspond to the version of the null specified in (3.1).
Theorem 4.2 establishes conditions under which the test $\phi_{n}$ is consistent (that is, has power approaching one for any distribution in the set of alternatives) whenever the selection rule $\rho$ is improvement-convergent.
Theorem 4.2.
[…]
$\lim_{n\rightarrow\infty}E_{P}\left[\phi_{n}\left(a_{0},\hat{a}^{\rho}_{1n}\right)\right]=1~{}$
We note here the technical condition in Theorem 4.2 which states that the selection rule $\rho$ should not select a candidate algorithm that is arbitrarily fair, even asymptotically. This is related to our discussion at the end of Section 3.2: if the selection rule selects an algorithm which converges to perfect fairness, then our bootstrap procedure is no longer consistent, and our proof strategy does not establish consistency of our test in this case. However, we emphasize that bootstrap consistency is not a necessary condition for our test to have reasonable power, and we demonstrate via simulation in Appendix D that our test displays favorable power even when employing candidate algorithms which are perfectly fair.
We expect that, for many datasets, it will be possible to reject the null using a naive selection rule that does not satisfy improvement convergence, as is the case in our empirical illustration in Section 5. However, Theorem 4.2 guarantees that the null will be asymptotically rejected (in cases where it should be) when the selection rule satisfies this property. In Appendix E we propose selection rules (constructed using mixed integer linear programs) for certain common utility specifications, which we conjecture to be improvement-convergent when $\mathcal{A}$ is the set of linear classifiers.
5. Empirical Application
We conclude by illustrating our approach in the setting of Obermeyer et al. (2019). The status quo algorithm is a commercial healthcare algorithm used by a large academic hospital to identify patients to target for a high-risk care management program. The algorithm assigns each patient a risk score, and identifies those patients with risk scores above the 97th percentile for automatic enrollment in the program. We use our proposed approach to evaluate the improvability of the hospital’s algorithm, finding that it is strictly FA-dominated within the class of linear classifiers. That is, it is possible to simultaneously strictly improve on the accuracy and the fairness of the algorithm. We further quantify the extent of possible improvement by testing the null hypotheses that the hospital’s algorithm is not $\delta$-fairness improvable and $\delta$-accuracy improvable for different values of $\delta$. We find that large improvements in fairness are possible without compromising on accuracy, while the reverse is not true.
5.1. Data and Classification Problem
The data includes observations from 48,784 patients, among which 43,202 self-report as White and 5,582 self-report as Black. Following Obermeyer et al. (2019), we take these to be the two group identities, denoted $g\in\{w,b\}$. Each patient $i$’s covariate vector $X_{i}\in\mathcal{X}$ includes 8 demographic variables (e.g., age and gender), 34 comorbidity variables (indicators of specific chronic illnesses), 13 cost variables (claimed cost broken down by type of cost), and 94 biomarker and medication variables. Although group identity $g$ is known for each patient in the dataset, the status quo algorithm omits this variable for prediction, so we do as well. Finally, the data set reports each patient $i$’s total number of active chronic illnesses in the subsequent year, which is interpreted as a measure of the patient’s true health needs.121212 In principle, enrollment into the high-risk care management program may affect the outcome variable $Y$. Although we expect that a single year of enrollment in this program will have only minor effects on the patients’ number of chronic illnesses (which $Y$ measures), we conduct a robustness check in Appendix G.2 in which we repeat our main analysis using the total number of active chronic illnesses in the same year as the enrollment decision. All results are qualitatively similar. We take this to be the patient’s type $Y_{i}\in\mathcal{Y}$.
An algorithm assigns each patient a decision $D_{i}\in\{0,1\}$ based on the covariate vector $X_{i}$, where 1 corresponds to an automatic enrollment into the care management program. We respect the capacity constraint of the status quo algorithm by restricting attention to algorithms that select 3% of the population for automatic enrollment.
Following Obermeyer et al. (2019), we evaluate both accuracy and fairness using the calibration utility function introduced in Example 2, i.e.
(5.1) $U^{g}(a)=E[Y\mid a(X)=1,G=g]$
and
$|U^{b}(a)-U^{w}(a)|=|E[Y\mid a(X)=1,G=b]-E[Y\mid a(X)=1,G=w]|.$
That is, an algorithm is more accurate if the expected number of health conditions is higher among both Black and White patients assigned to the program. It is more fair if it reduces the disparity in the expected number of health conditions among Black and White patients assigned to the program. [^13] Here and throughout, we drop the notational distinction between $U_{A}^{g}$ and $U_{F}^{g}$, as they are identical in this example.
5.2. Results
We implement our proposed approach in Section 3 to test the fairness- and accuracy- improvability of the hospital’s algorithm. We use $K=5$ iterations and select a $\beta=1/2$ fraction of the data to use for training. The status quo algorithm assigns each patient a risk score, and chooses those patients with risk scores in the top 3% for automatic enrollment. Our candidate algorithms similarly assign risk scores to patients, and enroll those patients with the highest risk scores, but use different risk score assignments compared to the status quo algorithm.
Specifically, to identify candidate algorithms, we first use the training data to compute an alternative risk score assignment rule $\hat{f}:\mathcal{X}\rightarrow\mathcal{Y}$, which maps each covariate vector into a predicted number of health conditions. Then we use the trained prediction rule $\hat{f}$ to rank all patients in the test data, and select the top 3% for automatic enrollment ($D=1$).141414 Obermeyer et al. (2019) find that the hospital’s algorithm was instead trained to predict a different outcome variable, namely health costs.
We consider several different methods for selecting a function $\hat{f}$—specifically, as the output of a linear regression, a lasso regression, or a random forest algorithm—but find that all of these yield similar results. [^15] The regularization parameter for LASSO is determined using 5-fold cross-validation. For random forest, we utilize a configuration of 300 trees with other hyperparameters set to their default values as specified in scikit-learn version 1.4.1. Figure 2 plots the average value (across the iterations of our procedure) of the utilities $U^{w}$ and $U^{b}$ under the status quo algorithm, as well as the average values for these utilities given candidate algorithms identified by each of the three methods mentioned above. This figure suggests that the candidate algorithms succeed in improving not only accuracy (each of $U^{w}$ and $U^{b}$ increase) but also fairness ($|U^{w}-U^{b}|$ decreases).
Figure 2. We report the average value of $U^{g}$ (i.e., number of active chronic diseases conditional on automatic enrollment) for each group $g$ across the $K=5$ iterations of our procedure. The algorithm is more accurate for group $g$ when $U^{g}$ is larger; so, moving towards the upper-right quadrant corresponds to improvements in accuracy. The algorithm is more fair when its $(U^{b},U^{w})$ pair is closer to the 45-degree line (corresponding to more balanced accuracy across the patient groups). This figure suggests that the candidate algorithms based on linear regression, LASSO, and random forest improve on both fairness and accuracy over the status quo algorithm used by the hospital.
We subsequently report findings for the linear regression model only, deferring the (nearly identical) results for the other models to Appendix G.1. Table 1 complements Figure 2 by reporting the $p$-values that emerge from our statistical test of the null hypothesis that the status quo algorithm is not strictly FA-dominated. Across the iterations of our procedure, the median $p$-value is smaller than $0.001$; thus, we reject the null at the 1% significance level. [^16] Recall that we reject at a $1\%$ significance level if the median $p$-value is less than $0.005$. Were the hospital to defend the disparate impact of its algorithm on account of the accuracy goal in (5.1), a federal funding agency could reject this business necessity claim with strong statistical guarantees. [^17] Our finding complements Obermeyer et al. (2019)’s result that training an algorithm to predict health needs (instead of health costs) leads to a more equal proportion of Black and White patients among those who are identified for automatic enrollment.
Accuracy (Black) Accuracy (White) Unfairness
$a_{1}$ $a_{0}$ $p_{b}$ $a_{1}$ $a_{0}$ $p_{w}$ $a_{1}$ $a_{0}$ $p_{f}$ $p$
Iteration 1 7.27 6.33 0.0000 7.24 5.15 0.0000 0.28 1.17 0.0000 0.0000
Iteration 2 7.37 6.37 0.0001 7.22 5.14 0.0000 0.15 1.23 0.0000 0.0001
Iteration 3 7.70 6.67 0.0000 7.24 5.17 0.0000 0.47 1.45 0.0000 0.0000
Iteration 4 7.42 6.38 0.0000 7.29 5.13 0.0000 0.13 1.25 0.0000 0.0000
Iteration 5 7.65 6.88 0.0017 7.32 5.28 0.0000 0.33 1.60 0.0000 0.0017
Table 1. The candidate algorithm $a_{1}$ in the table is based on linear regression. Reported $p$-values are computed via bootstrap with 10,000 iterations. The median $p$-value $p_{\mathrm{med}}$ is smaller than 0.001.
We further explore the tradeoff between improvements in accuracy and fairness by testing for $\delta$-fairness improvability, where we allow $\delta$ to vary. That is, we ask whether it is possible to improve on fairness by at least a factor of $\delta$ without compromising on accuracy. We find that we can reject the null at the usual 5% significance level for all $\delta\leq 0.68$. This means that even under an extremely stringent notion of what it means to improve on disparate impact—namely, that the status quo algorithm’s disparate impact is justified unless it is possible to halve disparate impact without compromising accuracy—the federal funding agency would still be able to conclude that improvements are feasible.
Figure 3. We test the null hypothesis that the status quo algorithm is not $\delta$-fairness improvable for $\delta\in[0,0.75]$, and report the corresponding median $p$-values. The null hypothesis is rejected for all $\delta\leq 0.68$.
We also test $\delta$-accuracy improvability. Although it is not directly relevant for the legal framework described in Section 1.1, it does provide a complementary angle on the fairness-accuracy tradeoffs in this application. We find that we can only conclude that joint improvements in accuracy of up to 7.5% are possible. We subsequently separately test for $\delta$-accuracy improvability for either group, and find that the binding dimension is accuracy improvement for Black patients. That is, while large gains in accuracy—up to 34%—can be achieved for White patients (without compromising accuracy for Black patients or fairness), we cannot conclude the same for Black patients.
(a) Test of $\delta$-joint accuracy improvability
(b) Test of $\delta$-accuracy improvability for White patients
(c) Test of $\delta$-accuracy improvability for Black patients
Figure 4. (A) We test the null hypothesis that the status quo algorithm is not $\delta$-jointly accuracy improvable, and reject for all $\delta\leq 0.075$. (B) We test the null hypothesis that the status quo algorithm is not $\delta$-accuracy improvable for White patients, and reject for all $\delta\leq 0.34$. (C) We test the null hypothesis that the status quo algorithm is not $\delta$-accuracy improvable for Black patients, and reject for all $\delta\leq 0.075$.
This finding can be explained in either of two ways. One possibility is that large improvements in accuracy for Black patients necessarily conflict with one of the two other criteria (i.e., it implies either a reduction in accuracy for White patients, or a reduction of fairness), while large improvements in accuracy for White patients can be achieved without sacrificing these other goals. Alternatively, it could be that the candidate algorithm simply does not substantially improve accuracy for Black patients. [^18] This could be, for example, because the measured covariates are not very predictive for this group. Our findings in Figure 4 are explained by the latter: At the threshold value of $\delta$ where the median $p$-value equals 0.025 (and we lose statistical significance), the binding $p$-value in the joint test is the one corresponding to a $\delta$-increase in accuracy for Black patients.
This is not a property of the data, but rather of our analysis in Figure 4. Since we fix the selection of candidate algorithms in this analysis, only the $p$-value pertaining to accuracy for Black patients varies with $\delta$ in panel (C). Thus for $\delta$ sufficiently large, this part of the test must be binding. (A similar statement holds for our other analyses in Figures 3 and 4.) To evaluate the inherent conflicts in accuracy and fairness more directly, one might instead allow the choice of candidate algorithm to depend on $\delta$, in which case all three $p$-values in our joint test would vary with $\delta$ and could in principle be the binding constraint. We leave such explorations to future work.
6. Conclusion
When a commercial algorithm has a disparate impact, it is important to understand whether it is possible to reduce that disparate impact without compromising on other business-relevant criteria, or if this level of disparate impact is necessitated by those other goals. We have designed a statistical approach to assess this, with three practical objectives in mind. First, since the appropriate definitions of disparate impact and business-relevant criteria vary substantially across applications, we want our framework to be flexible enough to accommodate any such definitions that may emerge in practice. Second, since there are often exogenous constraints on the algorithm space, we want our procedure to apply universally across algorithm classes. Finally, we would like for the approach to be transparent and simple to use for practitioners. The statistical framework that we propose delivers on these three counts.
The main drawback of our approach is its dependence on a selection rule for choosing a candidate algorithm. When we are able to reject the null, as in our application, then the optimality properties of the selection rule (in particular, whether it satisfies improvement convergence as defined in Section 4) do not matter. That is, no matter how naive or heuristic our selection rule is, we can conclude that the status quo algorithm is improvable. But when our test does not allow us to reject the null, there is in general an ambiguity: It may be that there is not sufficient evidence in the data to conclude existence of an improving algorithm; or, alternatively, it may be that the selection rule is not powerful enough to find this algorithm. In the latter case, re-running the same test with a different selection rule could lead us to reject the null. This ambiguity is resolved asymptotically when the selection rule is improvement-convergent, since Theorem 4.2 establishes that our proposed test is consistent under this condition. One interesting direction for future work is thus to provide sufficient conditions for a selection rule to be improvement-convergent in different relevant applications.
Appendix A Bootstrap algorithm
Let $\hat{P}_{n}$ be the empirical distribution of the data $\{(X_{i},Y_{i},G_{i})\}_{i=1}^{\ell_{n}}$ and let $\{(X^{*}_{i},Y^{*}_{i},G^{*}_{i})\}_{i=1}^{\ell_{n}}$ denote a bootstrap sample drawn i.i.d from $\{(X_{i},Y_{i},G_{i})\}_{i=1}^{\ell_{n}}$. Let $\hat{\theta}^{*}$, $\hat{A}^{*}_{tg}$, $\hat{F}^{*}_{tg}$ denote the bootstrap analogs of $\hat{\theta}$, $\hat{A}_{tg}$, $\hat{F}_{tg}$ computed using the bootstrap data. The bootstrap analogs of the test-statistics are given by
$\hat{T}_{g,n}^{*}=\hat{A}^{*}_{1g}-(1+\Delta_{g})\hat{A}_{0g}^{*}\text{ for $g\in\{r,b\}$}~{},$
$\hat{T}_{f,n}^{*}=\big{|}\hat{F}_{1r}^{*}-\hat{F}_{1b}^{*}\big{|}-(1-\Delta_{f})\big{|}\hat{F}_{0r}^{*}-\hat{F}_{0b}^{*}\big{|}~{}.$
The critical values for the test are based on the quantiles of the bootstrap distribution of the test statistics. Specifically, let $\Psi_{g,n}$ be the (conditional) cumulative distribution function of $\sqrt{\ell_{n}}(\hat{T}_{g,n}^{*}-\hat{T}_{g,n})$ given $\hat{P}_{n}$ and $\Psi_{f,n}$ be the (conditional) cumulative distribution function of $\sqrt{\ell_{n}}(\hat{T}_{f,n}^{*}-\hat{T}_{f,n})$ given $\hat{P}_{n}$. In practice these are not known exactly, but can be approximated via Monte-Carlo by taking $Q$ independent bootstrap samples from $\hat{P}_{n}$ and computing the empirical distribution
$\frac{1}{Q}\sum_{q=1}^{Q}I\{\sqrt{\ell_{n}}(\hat{T}^{*}_{s,n,q}-\hat{T}_{s,n})\leq x\}~{},$
for $s\in\{r,b,f\}$, where $\hat{T}^{*}_{s,n,q}$ denotes the bootstrap test statistic computed on the $q$th sample.
Appendix B Preliminary Results
Below we establish a sequence of preliminary results, which will be used to prove our Theorems 4.1 and 4.2. Section B.1 proves a result that allows us to reduce asymptotic validity of our joint test to asymptotic validity of tests of the component nulls. Section B.2 proves bootstrap consistency in a simpler environment where the sequence of candidate algorithms $a_{1n}$ is deterministic.
B.1. Intersection-Union Testing
Proposition B.1.
Let $\Omega$ denote a set of distributions for which $P\in\Omega$. Let $\Omega_{0}\subset\Omega$ denote the subset of distributions for which $H_{0}$ holds. Let $\phi^{(r)}_{n}\in\{0,1\}$, $\phi^{(b)}_{n}\in\{0,1\}$, $\phi^{(f)}_{n}\in\{0,1\}$ be tests of the component null hypotheses for accuracy and fairness as defined in Section 3.2. Let $\Omega_{0s}\subset\Omega$ for $s\in\{r,b,f\}$ denote the subsets of distributions for which each component null hypothesis $H_{0s}$ holds. Define the test $\phi_{n}=\phi^{(r)}_{n}\cdot\phi^{(b)}_{n}\cdot\phi^{(f)}_{n}$ which rejects if and only if each test $\phi^{(s)}_{n}$ rejects for $s\in\{r,b,f\}$.
  1. (1)
    If $\phi^{(s)}_{n}$ for $s\in\{r,b,f\}$ are asymptotically of level $\alpha$, i.e., if
    $\limsup_{n\rightarrow\infty}E_{P}[\phi_{n}^{(s)}]\leq\alpha~{},$
    for every $P\in\Omega_{0s}$, then $\phi_{n}$ is asymptotically of level $\alpha$, i.e.
    $\limsup_{n\rightarrow\infty}E_{P}[\phi_{n}]\leq\alpha~{},$
    for every $P\in\Omega_{0}$.
  2. (2)
    If $\phi^{(s)}_{n}$ for $s\in\{r,b,f\}$ are uniformly asymptotically of level $\alpha$, i.e. if
    $\limsup_{n\rightarrow\infty}\sup_{P\in\Omega_{0s}}E_{P}[\phi_{n}^{(s)}]\leq\alpha~{},$
    then $\phi_{n}$ is uniformly asymptotically of level $\alpha$, i.e.
    $\limsup_{n\rightarrow\infty}\sup_{P\in\Omega_{0}}E_{P}[\phi_{n}]\leq\alpha~{}.$
    Moreover, if there exists some distribution $P^{*}\in\Omega_{0}$ such that $E_{P^{*}}[\phi_{n}^{(s)}]\rightarrow\alpha$ for some $s\in\{r,b,f\}$, and $E_{P^{*}}[\phi_{n}^{(s^{\prime})}]\rightarrow 1$ for $s^{\prime}\neq s$, then
    $\lim_{n\rightarrow\infty}\sup_{P\in\Omega_{0}}E_{P}[\phi_{n}]=\alpha~{}.$
Proof.
To prove the first claim, note that by the definition of $\phi_{n}$, for any $P\in\Omega_{0}$,
$E_{P}[\phi_{n}]\leq E_{P}[\phi^{(s)}_{n}]~{},$
for every $s\in\{r,b,f\}$. By the definition of $\Omega_{0}$, $P\in\Omega_{0s}$ for some $s\in\{r,b,f\}$. Since $\phi^{(s)}_{n}$ is asymptotically of level $\alpha$, it follows that
$\limsup_{n\rightarrow\infty}E_{P}[\phi_{n}]\leq\limsup_{n\rightarrow\infty}E_{P}[\phi^{(s)}_{n}]\leq\alpha~{},$
as desired.
To prove the second claim, first note that by following the argument above, for each $P\in\Omega_{0}$ we obtain
$E_{P}[\phi_{n}]\leq\sup_{P^{\prime}\in\Omega_{0s}}E_{P^{\prime}}[\phi^{(s)}_{n}]~{},$
for some $s\in\{r,b,f\}$. It thus follows that
$\sup_{P\in\Omega_{0}}E_{P}[\phi_{n}]\leq\max\left\{\sup_{P^{\prime}\in\Omega_{0s}}E_{P^{\prime}}[\phi^{(s)}_{n}]:s\in\{r,b,f\}\right\}~{}.$
Since each $\phi^{(s)}_{n}$ is uniformly asymptotically of level $\alpha$, we then obtain that
(B.1) $\limsup_{n\rightarrow\infty}\sup_{P\in\Omega_{0}}E_{P}[\phi_{n}]\leq\alpha~{}.$
Next, note that from the definition of $\phi_{n}$,
$\sup_{P\in\Omega_{0}}E_{P}[\phi_{n}]$ $\geq E_{P^{*}}[\phi_{n}]$
$\geq\sum_{s}E_{P^{*}}[\phi_{n}^{(s)}]-2~{}.$
It thus follows by our assumptions on $P^{*}$ that
(B.2) $\liminf_{n\rightarrow\infty}\sup_{P\in\Omega_{0}}E_{P}[\phi_{n}]\geq\alpha~{}.$
Combining (B.1) and (B.2) we thus obtain that
$\lim_{n\rightarrow\infty}\sup_{P\in\Omega_{0}}E_{P}[\phi_{n}]=\alpha~{},$
as desired. ∎
B.2. Bootstrap consistency
To demonstrate Theorems 4.1 and 4.2, we first show that the bootstrap procedure defined in Appendix A is consistent when the test statistics are applied to an arbitrary deterministic sequence of candidate algorithms $a_{1n}\in\mathcal{A}$, i.e. that
[…]
as $n\rightarrow\infty$, for $s\in\{r,b,f\}$, where $d_{\infty}(\cdot,\cdot)$ denotes the Kolmogorov metric and $\mathcal{L}\{Z\}$ denotes the distribution function of a random variable $Z$. Note here that we have been careful to index each term in the expression by the total sample size $n$, since in principle each of these objects could vary with the candidate algorithm $a_{1n}$, but we suppress this dependence subsequently whenever we think it would not lead to confusion.
We prove bootstrap consistency in three steps. In the first step (Proposition B.2), we show that the asymptotic linear representations of the test statistics are asymptotically normal. In the second step (Proposition B.4), we demonstrate consistency of the bootstrap distribution for the asymptotic linear representations by invoking Theorem 1 of Mammen (1992). Finally in the third step we show that the bootstrap remains consistent when we move from the asymptotic linear representations to the original test statistics (Proposition B.5). Throughout this section we maintain Assumptions 1, 2 and 3.
B.2.1. Asymptotic normality of the test statistic
We first write the test statistics as sample averages. To do this, we define
$\Omega^{A}_{i}$ $=(u^{A}_{1r}(Z_{i},\theta),u^{A}_{0r}(Z_{i},\theta),u^{A}_{1b}(Z_{i},\theta),u^{A}_{0b}(Z_{i},\theta))~{},$
$\hat{\Omega}^{A}_{i}$ $=(u^{A}_{1r}(Z_{i},\hat{\theta}),u^{A}_{0r}(Z_{i},\hat{\theta}),u^{A}_{1b}(Z_{i},\hat{\theta}),u^{A}_{0b}(Z_{i},\hat{\theta}))~{},$
$\Omega^{F}_{i}$ $=(u^{F}_{1r}(Z_{i},\theta),u^{F}_{0r}(Z_{i},\theta),u^{F}_{1b}(Z_{i},\theta),u^{F}_{0b}(Z_{i},\theta))~{},$
$\hat{\Omega}^{F}_{i}$ $=(u^{F}_{1r}(Z_{i},\hat{\theta}),u^{F}_{0r}(Z_{i},\hat{\theta}),u^{F}_{1b}(Z_{i},\hat{\theta}),u^{F}_{0b}(Z_{i},\hat{\theta}))~{},$
$\mu_{i}$ $=(h_{i},\Omega^{A}_{i},\Omega^{F}_{i})~{},$
$\hat{\mu}_{i}$ $=(h_{i},\hat{\Omega}^{A}_{i},\hat{\Omega}^{F}_{i})~{},$
(B.3) $\Sigma_{n}$ $=E\left[(\mu_{i}-E[\mu_{i}])(\mu_{i}-E[\mu_{i}])^{T}\right]$
and
$\alpha_{r}$ $=(0_{J},1,-(1+\Delta_{r}),0_{6})~{},$
$\alpha_{b}$ $=(0_{J+2},1,-(1+\Delta_{b}),0_{4})~{},$
$\alpha_{f}$ $=\begin{pmatrix}0_{(J+4)\times 4}\\ I_{4}\\ \end{pmatrix}~{},$
where $0_{m}$ refers to a $m\times 1$ dimensional vector of 00s, $0_{m\times n}$ refers to a $m\times n$ dimensional matrix of 00s, and $I_{m}$ is the $m\times m$ dimensional identity matrix. We also define
$T_{r,n}$ $=E\left[\alpha_{r}^{T}\mu_{i}\right]~{},$
$T_{b,n}$ $=E\left[\alpha_{b}^{T}\mu_{i}\right]~{},$
$T_{f,n}$ $=\varphi(E\left[\alpha_{f}^{T}\mu_{i}\right])~{},$
where $\varphi(a,b,c,d)=|a-c|-(1-\Delta_{f})|b-d|$ and
$\hat{T}_{r,n}$ $=\frac{1}{\ell_{n}}\sum_{i=1}^{\ell_{n}}\alpha_{r}^{T}\hat{\mu}_{i}~{},$
$\hat{T}_{b,n}$ $=\frac{1}{\ell_{n}}\sum_{i=1}^{\ell_{n}}\alpha_{b}^{T}\hat{\mu}_{i}~{},$
$\hat{T}_{f,n}$ $=\varphi\left(\frac{1}{\ell_{n}}\sum_{i=1}^{\ell_{n}}\alpha_{f}^{T}\hat{\mu}_{i}\right)~{}.$
To demonstrate asymptotic normality of the test statistics, we derive their asymptotic linear representations. To do this, we use the additional definitions
$A^{1}_{tg}=E[\nabla u^{A}_{tg}(Z_{i},\theta)]~{},$
$F^{1}_{tg}=E[\nabla u^{F}_{tg}(Z_{i},\theta)]~{}.$
where we note that these are $J\times 1$ dimensional vectors, and
$\tilde{\alpha}_{r}$ $=(A_{1r}^{1}-A_{0r}^{1}(1+\Delta_{r}),1,-(1+\Delta_{r}),0_{6})~{},$
$\tilde{\alpha}_{b}$ $=(A_{1b}^{1}-A_{0b}^{1}(1+\Delta_{b}),0_{2},1,-(1+\Delta_{b}),0_{4})~{},$
$\tilde{\alpha}_{f}$ $=\left(\text{sign}(F_{1r}-F_{1b})(F_{1r}^{1}-F_{1b}^{1})-(1-\Delta_{f})\text{sign}(F_{0r}-F_{0b})(F_{0r}^{1}-F_{0b}^{1}),0_{4},\right.$
$\hskip 28.45274pt\left.\text{sign}(F_{1r}-F_{1b}),-(1-\Delta_{f})\text{sign}(F_{0r}-F_{0b}),-\text{sign}(F_{1r}-F_{1b}),(1-\Delta_{f})\text{sign}(F_{0r}-F_{0b})\right)~{}.$
Using these additional constructions, we can define the asymptotic linear representations of the test statistics and demonstrate their asymptotic normality. They are
$\tilde{T}_{r,n}$ $=T_{r,n}+\frac{1}{\ell_{n}}\sum_{i=1}^{\ell_{n}}\tilde{\alpha}_{r}^{T}(\mu_{i}-E[\mu_{i}]),$
$\tilde{T}_{b,n}$ $=T_{b,n}+\frac{1}{\ell_{n}}\sum_{i=1}^{\ell_{n}}\tilde{\alpha}_{b}^{T}(\mu_{i}-E[\mu_{i}]),$
$\tilde{T}_{f,n}$ $=T_{f,n}+\frac{1}{\ell_{n}}\sum_{i=1}^{\ell_{n}}\tilde{\alpha}_{f}^{T}(\mu_{i}-E[\mu_{i}]))~{}.$
First, we argue that the asymptotic linear representations are asymptotically normal.
Proposition B.2.
For $g\in\{r,b\}$,
$d_{\infty}\left(\mathcal{L}\left\{\sqrt{\ell_{n}}(\tilde{T}_{g,n}-T_{g,n})\right\},\mathcal{N}(0,\sigma_{g,n}^{2})\right)\rightarrow 0~{},$
and
$d_{\infty}\left(\mathcal{L}\left\{\sqrt{\ell_{n}}(\tilde{T}_{f,n}-T_{f,n})\right\},\mathcal{N}(0,\sigma_{f,n}^{2})\right)\rightarrow 0~{},$
where $\sigma_{g,n}^{2}=\tilde{\alpha}_{g}^{T}\Sigma_{n}\tilde{\alpha}_{g}$ and $\sigma_{f,n}^{2}=\tilde{\alpha}_{f}^{T}\Sigma_{n}\tilde{\alpha}_{f}$ for $g\in\{r,b\}$, and $d_{\infty}(\cdot,\cdot)$ is the Kolmogorov metric.
Proof.
For $g\in\{r,b\}$
$\sqrt{\ell_{n}}(\tilde{T}_{g,n}-T_{g,n})=\frac{1}{\sqrt{\ell_{n}}}\sum_{i}\tilde{\alpha}_{g}^{T}(\mu_{i}-E[\mu_{i}])~{}.$
Since $\mu_{i}$ is independent across $i$, the Lindeberg-Feller CLT implies that
[…]
where $\sigma_{g,n}^{2}=\tilde{\alpha}_{g}^{T}\Sigma_{n}\tilde{\alpha}_{g}$. Note that $\sigma_{g,n}^{2}>0$ by Assumption 2. It thus follows by Polya’s Theorem (Theorem 11.2.9 in Lehmann and Romano, 2022) that
$d_{\infty}\left(\mathcal{L}\left\{\sqrt{\ell_{n}}(\tilde{T}_{g,n}-T_{g,n})\right\},\mathcal{N}(0,\sigma_{g,n}^{2})\right)\rightarrow 0~{}.$
Similarly,
$\sqrt{\ell_{n}}(\tilde{T}_{f,n}-T_{f,n})=\frac{1}{\sqrt{\ell_{n}}}\sum_{i}\tilde{\alpha}_{f}^{T}(\mu_{i}-E[\mu_{i}])~{},$
and
$d_{\infty}\left(\mathcal{L}\left\{\sqrt{\ell_{n}}(\tilde{T}_{f,n}-T_{f,n})\right\},\mathcal{N}(0,\sigma_{f,n}^{2})\right)\rightarrow 0~{},$
where $\sigma_{f,n}^{2}=\tilde{\alpha}_{f}^{T}\Sigma_{n}\tilde{\alpha}_{f}$. Note also that $\sigma_{f,n}^{2}>0$ by Assumption 2. ∎
Second, we argue that the difference between the test statistics and their asymptotic linear representation are asymptotically negligible.
Proposition B.3.
Suppose there exists some $\xi>0$ such that $|U_{F}^{r}(a_{1n})-U_{F}^{b}(a_{1n})|\geq\xi$ for every $a_{1n}$ in our sequence of candidate algorithms. Then for each $g\in\{r,b\}$,
[…]
Proof.
The result for the accuracy test statistics follows from
$\hat{T}_{g,n}$ $=\alpha_{g}^{T}\frac{1}{\ell_{n}}\sum_{i}\hat{\mu}_{i}=\alpha_{g}^{T}\frac{1}{\ell_{n}}\sum_{i}(\mu_{i}+\nabla\mu_{i}(\hat{\theta}-\theta))+o_{p}(\ell_{n}^{-1/2})$
$=\tilde{\alpha}^{T}_{g}\frac{1}{\ell_{n}}\sum_{i}(0_{K},A_{i},F_{i})+\tilde{\alpha}^{T}_{g,n}\frac{1}{\ell_{n}}\sum_{i}(h_{i}-\theta,0_{8})+o_{p}(\ell_{n}^{-1/2})$
$=\frac{1}{\ell_{n}}\sum_{i}\tilde{\alpha}_{g}^{T}(\mu_{i}-(\theta,0_{8}))+o_{p}(\ell_{n}^{-1/2})=\tilde{T}_{g,n}+o_{p}(\ell_{n}^{-1/2})~{},$
where the last equality is because $\tilde{\alpha}_{g}^{T}(\theta,0_{8})+\alpha^{T}_{g}E[\mu_{i}]=\tilde{\alpha}_{g}^{T}E[\mu_{i}]$.
The second equality follows from a Taylor expansion, Assumption 2, and the fact that $||\hat{\theta}-\theta||=O_{p}(\ell_{n}^{-1/2})$ by a standard concentration argument (e.g. Markov’s inequality). The third equality follows from a law of large numbers, $||\hat{\theta}-\theta||=O_{p}(\ell_{n}^{-1/2})$, and some additional algebra.
The result for the fairness test statistic follows from
$\hat{T}_{f,n}$ $=\varphi\left(\frac{1}{\ell_{n}}\sum_{i}\alpha_{f}^{T}\hat{\mu}_{i}\right)=\varphi(E[\alpha_{f}^{T}\mu_{i}])+\nabla\varphi(E[\alpha_{f}^{T}\mu_{i}])\left(\frac{1}{\ell_{n}}\sum_{i}\alpha_{f}^{T}\hat{\mu}_{i}-E[\alpha_{f}^{T}\mu_{i}]\right)+o_{p}(\ell_{n}^{-1/2})$
$=\varphi(E[\alpha_{f}^{T}\mu_{i}])+\nabla\varphi(E[\alpha_{f}^{T}\mu_{i}])\alpha_{f}^{T}\left(\frac{1}{\ell_{n}}\sum_{i}(\mu_{i}-E[\mu_{i}])+E[\nabla\mu_{i}](\hat{\theta}-\theta)\right)+o_{p}(\ell_{n}^{-1/2})$
$=\varphi(E[\alpha_{f}^{T}\mu_{i}])+\nabla\varphi(E[\alpha_{f}^{T}\mu_{i}])\alpha_{f}^{T}(I_{K+8}+(E[\nabla\mu_{i}],0_{(K+8)\times 8}))\left(\frac{1}{\ell_{n}}\sum_{i}(\mu_{i}-E[\mu_{i}])\right)+o_{p}(\ell_{n}^{-1/2})$
$=T_{f,n}+\frac{1}{\ell_{n}}\sum_{i}\tilde{\alpha}_{f}^{T}(\mu_{i}-E[\mu_{i}])+o_{p}(\ell_{n}^{-1/2})=\tilde{T}_{f,n}+o_{p}(\ell_{n}^{-1/2})~{},$
where $I_{J+8}$ is a $(J+8)\times(J+8)$ dimensional identity matrix and $0_{(J+8)\times 8}$ is a $(J+8)\times 8$ dimensional matrix of 00s. Existence and continuity of $\nabla\varphi(\cdot)$ follows by Assumption 1 and the assumption that $|U_{F}^{r}(a_{1n})-U_{F}^{b}(a_{1n})|\geq\xi$.
The second equality follows from a Taylor expansion and because $\frac{1}{\ell_{n}}\sum_{i}\alpha_{f}^{T}\left(\hat{\mu}_{i}-E[\mu_{i}]\right)=O_{p}(\ell_{n}^{-1/2})$. The third equality is because $\frac{1}{\ell_{n}}\sum_{i}\left(\nabla\mu_{i}-E[\nabla\mu_{i}]\right)=O_{p}(\ell_{n}^{-1/2})$ and $||\hat{\theta}-\theta||=O_{p}(\ell_{n}^{-1/2})$. They both follow from Assumption 1 and a standard concentration argument.
The second-to-last equality is because $T_{f,n}=\varphi(E[\alpha_{f}^{T}\mu_{i}])$ and $\tilde{\alpha}_{f}^{T}=\nabla\varphi(E[\alpha_{f}^{T}\mu_{i}])\alpha_{f}^{T}(I_{J+8}+(E[\nabla\mu_{i}],0_{(J+8)\times 8}))$. This last result is because
$\nabla\varphi(E[\alpha^{T}_{f}\mu_{i}])=(\text{sign}(F_{1r}-F_{1b}),-(1-\Delta_{f})\text{sign}(F_{0r}-F_{0b}),-\text{sign}(F_{1r}-F_{1b}),-(1-\Delta_{f})\text{sign}(F_{0r}-F_{0b}))~{},$
$\alpha_{f}^{T}=\begin{pmatrix}0_{4\times(J+4)}&I_{4}\end{pmatrix}~{},\text{ and }$
$\left(I_{J+8}+(E[\nabla\mu_{i}],0_{(J+8)\times 8})\right)=\begin{pmatrix}I_{J}&0&0\\ A^{1}&I_{4}&0\\ F^{1}&0&I_{4}\\ \end{pmatrix}$
for $A^{1}=(A^{1}_{1r},A^{1}_{0r},A^{1}_{1b},A^{1}_{0b})$ and $F^{1}=(F^{1}_{1r},F^{1}_{0r},F^{1}_{1b},F^{1}_{0b})$ so that
$\nabla\varphi(E[\alpha_{f}^{T}\mu_{i}])\alpha_{f}^{T}(I_{J+8}+(E[\nabla\mu_{i}],0_{(J+8)\times 8}))$
$=\left(\text{sign}(F_{1r}-F_{1b})(F_{1r}^{1}-F_{1b}^{1})-(1-\Delta_{f})\text{sign}(F_{0r}-F_{0b})(F_{0r}^{1}-F_{0b}^{1}),0_{4},\right.$
$\hskip 28.45274pt\left.\text{sign}(F_{1r}-F_{1b}),-(1-\Delta_{f})\text{sign}(F_{0r}-F_{0b}),-\text{sign}(F_{1r}-F_{1b}),-(1-\Delta_{f})\text{sign}(F_{0r}-F_{0b})\right)$
$=\tilde{\alpha}_{f}~{}.$
It follows from Propositions B.2 and B.3 that for $g\in\{r,b\}$,
$d_{\infty}\left(\mathcal{L}\left\{\sqrt{\ell_{n}}(\hat{T}_{g,n}-T_{g,n})\right\},\mathcal{N}(0,\sigma_{g,n}^{2})\right)\rightarrow 0~{},$
and
$d_{\infty}\left(\mathcal{L}\left\{\sqrt{\ell_{n}}(\hat{T}_{f,n}-T_{f,n})\right\},\mathcal{N}(0,\sigma_{f,n}^{2})\right)\rightarrow 0~{}.$
B.2.2. Asymptotic normality of the bootstrap distribution
Let the (infeasible) bootstrap distributions of $\sqrt{\ell_{n}}(\tilde{T}_{g,n}-T_{g,n})$ and $\sqrt{\ell_{n}}(\tilde{T}_{f,n}-T_{f,n})$ be denoted by $\tilde{\Psi}_{g,n}$ and $\tilde{\Psi}_{f,n}$. We say that these distributions are infeasible because they depend on unknown population parameters.
Proposition B.4.
[…]
Proof.
This follows from Proposition B.2 above and Theorem 1 of Mammen (1992). ∎
Proposition B.5.
Suppose $|U_{F}^{r}(a_{1n})-U_{F}^{b}(a_{1n})|\geq\xi$ for some $\xi>0$, for every $a_{1n}$ in our sequence of candidate algorithms. For $s\in\{r,b,f\}$,
[…]
as $n\rightarrow\infty$.
Proof.
Let $P_{*}$ denote the bootstrap distribution conditional on the data. By applying the logic of Proposition B.3 to the bootstrap data, it follows that
$\sqrt{\ell_{n}}(\hat{T}^{*}_{s,n}-\tilde{T}^{*}_{s,n})=r^{*}_{s,n}~{},$
[…]
$P_{*}\left(|r^{*}_{s,n_{k}}|>\epsilon\right)\xrightarrow{a.s.}0~{},$
$\sqrt{\ell_{n}}(\hat{T}_{s,n_{k}}-\tilde{T}_{s,n_{k}})\xrightarrow{a.s.}0~{},$
$d_{\infty}(\tilde{\Psi}_{s,n_{k}},\mathcal{N}(0,\sigma^{2}_{s,n_{k}}))\xrightarrow{a.s.}0~{},$
$\sigma_{s,n_{k}}\rightarrow\sigma^{\dagger}>0~{},$
where the second convergence statement follows from Proposition B.3, the third from Proposition B.4, and the last convergence statement follows from Assumptions 2 and 3. We thus have that
[…]
conditional on the data, with probability one. It follows by Polya’s Theorem (Theorem 11.2.9 in Lehmann and Romano, 2022) that
$d_{\infty}\left(\Psi_{s,n_{k}},N(0,\sigma^{2}_{s,n_{k}})\right)\xrightarrow{a.s.}0~{},$
and thus by Propositions B.2 and B.3,
$d_{\infty}\left(\mathcal{L}\left\{\sqrt{\ell_{n_{k}}}(\hat{T}_{s,n_{k}}-T_{s,n_{k}})\right\},\Psi_{s,n_{k}}\right)\xrightarrow{a.s.}0~{}.$
Since we have established that for every subsequence there exists a further subsequence along which the convergence holds almost surely, the original sequence converges in probability, as desired. ∎
Appendix C Proofs of Main Results
C.1. Proof of Theorem 4.1
We will first show that for any deterministic sequence of algorithms $a_{1n}\in\mathcal{A}$, and any $P$ satisfying the null hypothesis,
(C.1) $\limsup_{n\rightarrow\infty}E_{P}[\phi_{n}(a_{0},a_{1n})]\leq\alpha~{}.$
By Proposition B.1 (1) it suffices to check that each test $\phi_{n}^{(s)}$ is asymptotically of level $\alpha$ for distributions $P$ contained in $\Omega_{0s}$, $s\in\{r,b,f\}$. We prove the result for the fairness test since the other results follow from similar arguments. By definition,
$E_{P}[\phi_{n}^{(f)}]$ $=P(\sqrt{\ell_{n}}\hat{T}_{f,n}<c^{*}_{\alpha})$
(C.2) $\leq P(\sqrt{\ell_{n}}(\hat{T}_{f,n}-T_{f,n})<c^{*}_{\alpha})~{},$
since $T_{f,n}\geq 0$ for $P\in\Omega_{0f}$.
Further note by Assumption 1 that for any $P\in\Omega_{0f}$,
$|U_{F}^{r}(a_{1n})-U_{F}^{b}(a_{1n})|\geq|F_{0r}-F_{0b}|(1-\Delta_{f})>0$
for every $a_{1n}$ in our sequence of candidate algorithms. Thus by Proposition B.5, for any $\epsilon>0$ and $n$ sufficiently large,
$P\left\{d_{\infty}\left(\mathcal{L}\left\{\sqrt{\ell_{n}}(\hat{T}_{f,n}-T_{f,n})\right\},\Psi_{f,n}\right)\leq\frac{\epsilon}{2}\right\}\geq 1-\frac{\epsilon}{2}~{}$
where $d_{\infty}(\cdot,\cdot)$ denotes the Kolmogorov metric and $\mathcal{L}\{Z\}$ denotes the distribution function of a random variable $Z$. Thus by Lemma A.1. (vii) in Romano and Shaikh (2012), we have that for $n$ sufficiently large,
$P(\sqrt{\ell_{n}}(\hat{T}_{f,n}-T_{f,n})\geq c^{*}_{\alpha})\geq 1-\alpha-\epsilon~{}.$
Combining this with (C.2), it follows that for $n$ sufficiently large,
$P(\sqrt{\ell_{n}}\hat{T}_{f,n}<c^{*}_{\alpha})\leq 1-(1-\alpha-\epsilon)=\alpha+\epsilon~{},$
from which we can conclude that
$\limsup_{n\rightarrow\infty}E_{P}[\phi_{n}^{(f)}]\leq\alpha~{}.$
Repeating this argument for the other two component nulls, we obtain the desired (C.1).
Next we argue that the result holds for the random sequence $\hat{a}^{\rho}_{1n}$. By the law of iterated expectations,
$E_{P}[\phi_{n}(a_{0},\hat{a}^{\rho}_{1n})]=E\left[E_{P}[\phi_{n}(a_{0},\hat{a}^{\rho}_{1n})\mid S^{n}_{train}]\right]~{}.$
By the independence of $S^{n}_{train}$ from $S^{n}_{test}$, $E_{P}[\phi_{n}(a_{0},\hat{a}^{\rho}_{1n})\mid S^{n}_{train}]=h(\hat{a}^{\rho}_{1n})$, where $h(a)=E_{P}[\phi_{n}(a_{0},a)]$. Thus by Fatou’s lemma and (C.1), we obtain
$\limsup_{n\rightarrow\infty}E_{P}[\phi_{n}(a_{0},\hat{a}^{\rho}_{1n})]\leq E\left[\limsup_{n\rightarrow\infty}h(\hat{a}^{\rho}_{1n})\right]\leq\alpha~{},$
as desired.
C.2. Proof of Corollary 4.1
By Markov’s inequality
$E_{P}[\psi_{n}]\leq\frac{2}{K}\sum_{k=1}^{K}E_{P}[\phi_{n,k}(a_{0},\hat{a}^{\rho}_{1nk})]~{}.$
The result then follows immediately from Theorem 4.1.
C.3. Proof of Theorem 4.2
Consider a deterministic sequence of algorithms $a_{1n}\in\mathcal{A}$ satisfying:
  1. (1)
    there exists a $\xi>0$ such that $|U_{F}^{r}(a_{1n})-U_{F}^{b}(a_{1n})|\geq\xi$ for each $a_{1n}$, and
  2. (2)
    $U^{g}_{h}(a_{1n})\rightarrow U^{g}_{h}(\gamma)$ for each $g\in\{r,b\}$ and $h\in\{A,F\}$, where $\gamma\in\mathcal{A}$ is an algorithm that $\delta$-fairness improves on the status quo algorithm $a_{0}$.
We first argue that $E_{P}[\phi_{n}^{(s)}]\rightarrow 1$ for $s\in\{r,b,f\}$. We show the result for $\phi_{n}^{(g)}$ for $g\in\{r,b\}$, as the result for fairness follows similarly. We proceed along subsequences. By Assumption 3, for every subsequence there exists a further subsequence (which we index by $n_{k}$) such that $\sigma^{2}_{g,n_{k}}\rightarrow\sigma^{2}$ for some $\sigma^{2}>0$. Thus by Proposition B.5
[…]
Along a further subsequence (which we continue to index by $n_{k}$), we obtain that (C.3) holds almost surely, and thus by Lemma 11.2.1 in Lehmann and Romano (2022), $c^{*}_{1-\alpha}\xrightarrow{a.s}c_{1-\alpha}$ along this subsequence, where $c_{1-\alpha}$ denotes the $1-\alpha$ quantile of a $N(0,\sigma^{2})$ distribution. We thus obtain that, along this subsequence,
$E_{P}[\phi_{n_{k}}^{(g)}]$ $=P\left(\sqrt{\ell_{n_{k}}}\hat{T}_{g,n_{k}}>c^{*}_{1-\alpha}\right)$
$=P\left(\sqrt{\ell_{n_{k}}}\left(\hat{T}_{g,n_{k}}-T_{g,n_{k}}\right)>c^{*}_{1-\alpha}-\sqrt{\ell_{n_{k}}}T_{g,n_{k}}\right)\rightarrow 1~{},$
where the convergence follows from Corollary 11.3.1 in Lehmann and Romano (2022) and the fact that, since $\gamma$ is a fairness-improvement,
$c^{*}_{1-\alpha}-\sqrt{\ell_{n_{k}}}T_{g,n_{k}}\xrightarrow{a.s}-\infty~{}.$
Since we have established that for every subsequence there exists a further subsequence along which $E_{P}[\phi_{n_{k}}^{(g)}]\rightarrow 1$, we obtain $E_{P}[\phi_{n}^{(g)}]\rightarrow 1$. Similar arguments establish $E_{P}[\phi_{n}^{(f)}]\rightarrow 1$. Next, we argue that these results continue to hold when we replace the sequence $a_{1n}$ with $\hat{a}^{\rho}_{1n}$. To that end, once again we argue along subsequences. For each subsequence $\phi_{n_{k}}^{(g)}$, there exists a further subsequence such that $U^{g}_{h}(\hat{a}^{\rho}_{1n})\xrightarrow{a.s}U^{g}_{h}(\gamma)$ for each $g\in\{r,b\}$, $h\in\{A,F\}$. Then by the law of iterated expectations and the dominated convergence theorem,
$E_{P}[\phi_{n_{k}}^{(g)}]=E[E[\phi_{n_{k}}^{(g)}|S^{n}_{train}]]\rightarrow 1~{},$
so that the result follows along the original sequence. Finally we establish the main result of the theorem: by construction
$E_{P}[\phi_{n}(a_{0},\hat{a}^{\rho}_{1n})]\geq E_{P}[\phi_{n}^{(r)}]+E_{P}[\phi_{n}^{(b)}]+E_{P}[\phi_{n}^{(f)}]-2\rightarrow 1~{},$
as desired.
Appendix D Simulation Exercise
In this section we study the effect of employing arbitrarily fair candidate algorithms on the power of our bootstrap procedure for testing fairness. To focus on this specific feature of the test, we abstract away from the details of our procedure and explicitly generate values of the status-quo and candidate fairness measures using a data generating process given by
$\begin{bmatrix}\Gamma_{0}\\ \Gamma_{1}\end{bmatrix}\sim N\left(\begin{bmatrix}1.25\\ \eta\end{bmatrix},\begin{bmatrix}10&7.3\\ 7.3&11.91\end{bmatrix}\right)~{}.$
Here, $\Gamma_{0}$ and $\Gamma_{1}$ represent random draws of the utility discrepancy between groups (i.e. realizations of $u^{F}_{tr}(Z_{i},\theta)-u^{F}_{tb}(Z_{i},\theta)$) for the status-quo and candidate algorithms, respectively. Our null hypothesis of interest could thus be restated as
$H_{0}:\left|E[\Gamma_{1}]\right|-|E[\Gamma_{0}]|\geq 0~{},$
against the alternative
$H_{1}:|E[\Gamma_{1}]|-|E[\Gamma_{0}]|<0~{}.$
The mean of $\Gamma_{0}$ and the covariance matrix for $[\Gamma_{0},\Gamma_{1}]$ were calibrated using the test sample from Iteration 4 (the median split) in Table 1. Our objective is to study the power of the test $\phi_{n}^{(f)}$ for a range of values of the mean fairness of the candidate algorithm given by $\eta\in[0,1.75]$. For each $\eta$, we report the rejection probability of a $5\%$ level test based on $2,000$ bootstrap draws, computed using $5,000$ Monte Carlo iterations. To simplify the computation, we consider much smaller test set sample sizes ($\ell_{n}\in\{100,200,400\}$) than the sample size available in the empirical application, but we simultaneously reduce the amount of noise in the problem considerably, by re-normalizing the covariance matrix so that $\text{Var}(\Gamma_{0})=10$.
Figure 5 presents the computed power curves for each sample size. In all cases, we see that the test controls size when $\eta\geq 1.25$, as expected. As candidate fairness approaches zero, we see that power generally increases. However, when $\ell_{n}\in\{100,200\}$, we also observe that the power of the test dips slightly as we approach perfect fairness. We conjecture that this may be related to the fact that the critical value generated by our bootstrap procedure could be less well-behaved at points near perfect fairness, as mentioned in the discussion following Theorem 4.2. However, as expected given our theoretical results, this feature of the power curve disappears once the test sample is sufficiently large (in this case, $\ell_{n}=400$).
(a) Sample size $\ell_{n}=100$
(b) Sample size $\ell_{n}=200$
(c) Sample size $\ell_{n}=400$
Figure 5. Rejection probabilities for various levels of candidate fairness, for sample sizes $\ell_{n}\in\{100,200,400\}$.
Appendix E Search for an Improving Algorithm
In this section we present a collection of systematic methods for finding a candidate algorithm which may improve over the status quo for two specific utility specifications: the classification rate utility (Example 1) and the calibration utility (Example 2). Before proceeding, we note that our ultimate test for fairness/accuracy-improvability is valid regardless of the method used to construct the candidate algorithm, and good examples of possible methods are already available in the literature for some utility specifications (see Section 1.2.) With this in mind, our goal in this section is to provide methods which are specifically tailored to the problem of interest.
As in Section 2.2, let $a_{0}$ denote a status quo algorithm and let $\mathcal{A}$ denote some fixed class of algorithms. Our objective is to find a candidate algorithm $a_{1}\in\mathcal{A}$ with which we can test for a fairness and/or accuracy improvement relative to $a_{0}$, using the test outlined in Section 3.2. For a given class of algorithms $\mathcal{A}$, let $\mathcal{U}(\mathcal{A})$ denote the feasible set of expected accuracy and fairness utilities that can be realized by algorithms in $\mathcal{A}$. Consider first the search for an algorithm that $\delta$-fairness improves on $a_{0}$. The following optimization problem
(E.1) $\inf_{U^{r}_{F},U^{b}_{F},U^{r}_{A},U^{b}_{A}\in\mathbb{R}_{+}}$ $|U^{r}_{F}-U^{b}_{F}|$
s.t. $U^{r}_{A}>U^{r}_{A}(a_{0})$
$U^{b}_{A}>U^{b}_{A}(a_{0})$
$(U^{r}_{F},U^{b}_{F},U^{r}_{A},U^{b}_{A})\in\mathcal{U}(\mathcal{A})~{}$
returns the smallest possible unfairness, subject to feasibility and no reduction in accuracy. The status quo algorithm is $\delta$-fairness improvable if and only if the optimal value of this optimization problem is strictly less than $|U^{r}_{F}(a_{0})-U^{b}_{F}(a_{0})|(1-\delta)$. We could alternatively search for an algorithm that $\delta$-improves on accuracy for group (say, for instance) $r$ in the same way. Such an algorithm exists if and only if the optimal value of the following optimization problem is strictly greater than $U^{r}_{A}(a_{0})(1-\delta)$:
(E.2) $\sup_{U^{r}_{F},U^{b}_{F},U^{r}_{A},U^{b}_{A}\in\mathbb{R}_{+}}$ $U^{r}_{A}$
s.t. $U^{b}_{A}>U^{b}_{A}(a_{0})$
$|U^{r}_{F}-U^{b}_{F}|<|U^{r}_{F}(a_{0})-U^{b}_{F}(a_{0})|$
$(U^{r}_{F},U^{b}_{F},U^{r}_{A},U^{b}_{A})\in\mathcal{U}(\mathcal{A})~{}.$
We propose to search for candidates $a_{1}\in\mathcal{A}$ by solving feasible versions of the optimization problems (E.1), (E.2) given a set of training data $\{(X_{j},Y_{j},G_{j})\}_{j=1}^{m}$. Let $\hat{U}^{g}_{h}(a)$ denote the empirical utility of algorithm $a$ for group $g\in\{r,b\}$, $h\in\{A,F\}$ computed using the sample $\{(X_{j},Y_{j},G_{j})\}_{j=1}^{m}$ (for the status quo algorithm $a_{0}$ we will often denote this as $\hat{U}^{g}_{0h}$). Let
[…]
be a sample analog of the feasible set. A candidate for improvement could now be computed by solving the sample analog of optimization problem (E.1) given by
(E.3) $\min_{U^{r}_{F},U^{b}_{F},U^{r}_{A},U^{b}_{A}\in\mathbb{R}_{+}}$ $|U^{r}_{F}-U^{b}_{F}|$
s.t. $U^{r}_{A}\geq\hat{U}^{r}_{0A}+\iota$
$U^{b}_{A}\geq\hat{U}^{b}_{0A}+\iota$
$(U^{r}_{F},U^{b}_{F},U^{r}_{A},U^{b}_{A})\in\widehat{\mathcal{U}}(\mathcal{A})~{}$
where $\iota\geq 0$ is some small positive constant. Solving (E.3) may not be tractable in general. In the rest of this section we consider tractable versions of problem (E.3) for specific utility specifications, by reformulating (E.3) as a mixed integer linear program (MILP). Modern optimization libraries are capable of solving MILPs of moderate size (recent applications in econometrics include Florios and Skouras, 2008; Kitagawa and Tetenov, 2018; Chen and Lee, 2018; Mbakop and Tabord-Meehan, 2021).
E.1. Classification Utility
Consider a setting with $\mathcal{Y}=\mathcal{D}=\{0,1\}$, $w_{A}=w_{F}=w(x,y,d)=1$, and $u_{A}=u_{F}=u(x,y,d)=\mathbbm{1}(y=d)$, so that our objective is to find a classifier $a:\mathcal{X}\rightarrow\{0,1\}$. Moreover, we fix $\mathcal{A}$ to be the set of linear classifiers:
Definition E.1.
Algorithm $a:\mathbb{\mathcal{X}}\rightarrow\{0,1\}$ is a linear classifier if $a(x)=\mathbbm{1}(\beta^{T}x\geq 0)$ for some $\beta\in\mathbb{R}^{d}$. From now on we use $\mathcal{A}$ to denote the set of all linear classifiers. (We assume that each vector $X_{i}$ contains a $1$ in its first entry, so it is without loss to set the threshold to zero.)
Given these restrictions, we can operationalize (E.3) via the following mixed integer linear program (MILP) formulation of the feasible set $\widehat{\mathcal{U}}(\mathcal{A})$ (note here that we set the fairness and accuracy utilities to be equal, so that the feasible set consists of utility pairs):
(E.4) $\widehat{\mathcal{U}}(\mathcal{A})=\left\{(U^{r},U^{b})\in[0,1]^{2}\ \begin{tabular}[]{|l}$U^{r}=\frac{1}{n_{r}}\sum_{G_{j}=r}\left((1-Y_{j})+(2Y_{j}-1)D_{j}\right)$\\[4.0pt] $U^{b}=\frac{1}{n_{b}}\sum_{G_{j}=b}\left((1-Y_{j})+(2Y_{j}-1)D_{j}\right)$\\[4.0pt] $\frac{X_{j}^{T}\beta}{C_{j}}<D_{j}\leq 1+\frac{X_{j}^{T}\beta}{C_{j}}\,\,\forall\hskip 2.84526pt1\leq j\leq m,\text{for some $\beta\in\mathcal{B}$}$\\[4.0pt] $D_{j}\in\{0,1\}\,\,\forall\hskip 2.84526pt1\leq j\leq m$\end{tabular}\right\}$
where $n_{g}=\sum_{G_{j}=g}1$, and $C_{j}$ is chosen to satisfy $C_{j}>\sup_{\beta\in\mathcal{B}}|X_{j}^{T}\beta|$, with $\mathcal{B}$ some compact set such that we restrict $\beta\in\mathcal{B}$. The first, second and fourth conditions guarantee that there is binary vector of decisions $(D_{j})_{j=1}^{m}$ (where $D_{j}=1$ corresponds to classifying individual $j$ as $1$) yielding sample utilities $U^{r}$ and $U^{b}$ under classification loss. The inequalities in the third condition further constrain that these decisions are consistent with a linear classifier, since if $X_{j}^{T}\beta$ is weakly positive then the constraint $D_{j}\in\{0,1\}$ implies $D_{j}=1$, while if $X_{j}^{T}\beta$ is strictly negative then the constraint $D_{j}\in\{0,1\}$ implies $D_{j}=0$. We can then use this characterization of $\widehat{\mathcal{U}}(\mathcal{A})$ to construct the following MILP formulation of (E.3):
(E.5) $\min_{(t,U^{r},U^{b})\in[0,1]^{3}}$ $t$
s.t. $t\geq U^{r}-U^{b}$
$t\geq-(U^{r}-U^{b})$
$U^{r}\geq\hat{U}^{r}_{0}+\iota$
$U^{b}\geq\hat{U}^{b}_{0}+\iota$
$(U^{r},U^{b})\in\widehat{\mathcal{U}}(\mathcal{A})~{},$
where we note that the first two constraints reformulate the absolute value in the objective function as a set of linear inequalities. We could also operationalize a feasible version of problem (E.2) as a MILP as well:
(E.6) $\max_{(U^{r},U^{b})\in[0,1]^{2}}$ $U^{r}$
s.t. $U^{b}\geq\hat{U}^{b}_{0}+\iota$
$U^{r}-U^{b}\leq|\hat{U}^{r}_{0}-\hat{U}^{b}_{0}|-\iota$
$U^{r}-U^{b}\geq-|\hat{U}^{r}_{0}-\hat{U}^{b}_{0}|+\iota$
$(U^{r},U^{b})\in\widehat{\mathcal{U}}(\mathcal{A})~{}.$
Let $\beta_{f}$ denote a coefficient vector that achieves the minimum value of problem (E.5). Then a candidate algorithm for $\delta$-fairness improvement is $a_{f}(x)=\mathbbm{1}(\beta_{f}^{T}x\geq 0)$. Let $\beta_{r}$ denote a coefficient vector that achieves the minimum value of of problem (E.6). Then a candidate algorithm for an $\delta$accuracy improvement for group $r$ is $a_{r}(x)=\mathbbm{1}(\beta_{r}^{T}x\geq 0)$.
E.2. Calibration Utility
Consider a setting in Example 2: we have $\mathcal{Y}\subseteq\mathbb{R}$, $\mathcal{D}=\{0,1\}$,
$u(x,y,d)=y\mathbbm{1}(d=1),\quad w(x,y,d)=\mathbbm{1}(d=1),\quad u_{A}=u_{F}=u,\quad w_{A}=w_{F}=w.$
Our objective is to find a linear classifier. As in our empirical application, we set the capacity constraint by restricting to algorithms that select a fraction $\kappa$ of the population. With a slight abuse of notation, let $G_{i}\equiv 1$ when $G_{i}=b$; otherwise let $G_{i}\equiv 0$. The error pair $(U^{r},U^{b})\in\mathbb{R}_{+}^{2}$ is in the feasible set $\hat{\mathcal{U}}(\mathcal{A})$ if and only if there exist $\beta\in\mathcal{B}$ such that
(E.7) $\left\{\begin{matrix}\displaystyle{U^{b}=\left(\sum_{i}G_{i}D_{i}\right)^{-1}\sum_{i}Y_{i}G_{i}D_{i}}\\[8.0pt] \displaystyle{U^{r}=\left(\sum_{i}(1-G_{i})D_{i}\right)^{-1}\sum_{i}Y_{i}(1-G_{i})D_{i}}\\[8.0pt] \displaystyle{\sum_{i}D_{i}=\kappa m}\\[8.0pt] \frac{X_{i}^{\top}\beta}{C_{i}}<D_{i}\leq 1+\frac{X_{i}^{\top}\beta}{C_{i}}\ \forall i\\[8.0pt] D_{i}\in\{0,1\}\ \forall i\end{matrix}\right.$
where $m$ in the third condition denotes the size of the train set.
We can formulate mixed integer programming (MIP) corresponding to (E.5) and (E.6) for the current $\hat{\mathcal{U}}(\mathcal{A})$. However, the optimization problem is not linear in decision variables since the first and second conditions in (E.7) are nonlinear. Below, we show that we could operationalize the problem by transforming it into a MILP.
To describe the transformation, let us introduce some variables:
$z_{1i}\equiv Y_{i}G_{i},\ z_{0i}\equiv Y_{i}(1-G_{i}),\ z_{1}\equiv(z_{1i})_{i},\ z_{0}\equiv(z_{0i})_{i},\ d\equiv(D_{i})_{i},\ g\equiv(G_{i})_{i}.$
Also, let $\mathbbm{1}$ denote the vector whose elements are all $1$. Then MIP corresponding (E.5), say $(P)$, can be written as follows:
$\mathrm{(P)}\left[\begin{matrix}\displaystyle{\min_{\beta,d,U^{b},U^{r}}}&\displaystyle{\left|U^{b}-U^{r}\right|}\\[8.0pt] \mathrm{s.t.}&U^{b}=\displaystyle{\frac{z_{1}^{\top}d}{g^{\top}d}}\\[8.0pt] &U^{r}=\displaystyle{\frac{z_{0}^{\top}d}{(1-g)^{\top}d}}\\[8.0pt] &U^{b}\geq\hat{U}^{b}_{0}+\iota\\ &U^{r}\geq\hat{U}^{r}_{0}+\iota\\ &\displaystyle{\mathbbm{1}^{\top}d=\kappa m}\\ &\displaystyle{\frac{X_{i}^{\top}\beta}{C_{i}}}<D_{i}\leq 1+\displaystyle{\frac{X_{i}^{\top}\beta}{C_{i}}}\\[8.0pt] &D_{i}\in\{0,1\},\beta\in\mathcal{B}\end{matrix}\right.$
Note that for each group there exists at least one $i$ with $D_{i}=1$ at the optimal solution of $(P)$; otherwise the value of the objective function is $\infty$. Consider the following change of variables:
(E.8) $t_{1}\equiv\frac{1}{g^{\top}d},\quad t_{0}\equiv\frac{1}{(1-g)^{\top}d},\quad s_{1}\equiv t_{1}d,\quad s_{0}\equiv t_{0}d.$
Then we can formulate the following MILP that corresponds to the original problem:
$\mathrm{(P^{\prime})}\left[\begin{matrix}\displaystyle{\min_{\begin{subarray}{c}\beta,d,U^{b},U^{r},\\ s_{1},s_{0},t_{1},t_{0}\end{subarray}}}&\displaystyle{\left|U^{b}-U^{r}\right|}\\[8.0pt] \mathrm{s.t.}&U^{b}=z_{1}^{\top}s_{1}\\ &U^{r}=z_{0}^{\top}s_{0}\\ &U^{b}\geq\hat{U}^{b}_{0}+\iota\\ &U^{r}\geq\hat{U}^{r}_{0}+\iota\\ &\displaystyle{\mathbbm{1}^{\top}d=\kappa m}\\ &g^{\top}s_{1}=1\\ &(1-g)^{\top}s_{0}=1\\[8.0pt] &\displaystyle{\frac{X_{i}^{\top}\beta}{C_{i}}}<D_{i}\leq\displaystyle{\left(1+\frac{X_{i}^{\top}\beta}{C_{i}}\right)}\ \forall i\\[8.0pt] &s_{1i}\leq t_{1}\ \forall i\\ &s_{1i}\leq D_{i}\ \forall i\\ &s_{1i}\geq t_{1}-(1-D_{i})\ \forall i\\ &s_{0i}\leq t_{0}\ \forall i\\ &s_{0i}\leq D_{i}\ \forall i\\ &s_{0i}\geq t_{0}-(1-D_{i})\ \forall i\\ &t_{1}\in(0,1],t_{0}\in(0,1],s_{1i}\geq 0,s_{0i}\geq 0,U^{b}\geq 0,U^{r}\geq 0,\\ &D_{i}\in\{0,1\},\beta\in\mathcal{B}\end{matrix}\right.$
The following proposition guarantees that we can utilize the MILP $(P^{\prime})$ to find a candidate algorithm for calibration utilities. Therefore, if the problem size is moderate, we can find a candidate algorithm using modern optimization libraries.
Proposition E.1.
The optimal values of $(P)$ and $(P^{\prime})$ coincide. Moreover, the optimal solution to $(P)$ can be computed by using the optimal solution to $(P^{\prime})$.
Proof.
Denote the optimal values of $(P)$ and $(P^{\prime})$ by OPT$(P)$ and OPT$(P^{\prime})$, respectively. For each solution feasible in $(P)$, there is a corresponding solution feasible in $(P^{\prime})$. Specifically, given $(\beta,d,U^{r},U^{b})$ feasible in $(P)$, $(\beta,d,U^{r},U^{b},s_{1},s_{0},t_{1},t_{0})$, where $(s_{1},s_{0},t_{1},t_{0})$ are defined in (E.8), is feasible in $(P^{\prime})$ and gives the same value of the objective funtion. Thus, OPT$(P^{\prime})$leq OPT$(P)$.
To show OPT$(P^{\prime})$geq OPT$(P)$, we show that for each feasible solution in $(P^{\prime})$, there exists a corresponding solution feasible in $(P)$ with the same value of the objective function. It suffices to show that
$\frac{s_{1i}}{t_{1}}=\frac{s_{0i}}{t_{0}}=D_{i},$
for each $i$ if $(\beta,d,U^{r},U^{b},s_{1},s_{0},t_{1},t_{0})$ is feasible in $(P^{\prime})$, which guarantees that (E.8) indeed holds and $(\beta,d,U^{r},U^{b})$ is feasible in $(P)$. Below, we will show that the conditions
(E.9) $s_{1i}$ $\leq t_{1},$
(E.10) $s_{1i}$ $\leq D_{i},$
(E.11) $s_{1i}$ $\geq t_{1}-(1-D_{i}),$
(E.12) $s_{1i}$ $\geq 0,\text{ and }$
(E.13) $t_{1}$ $\in(0,1],$
imply $s_{1i}/t_{1}=D_{i}$ ($s_{0i}/t_{0}=D_{i}$ can be shown in the same manner). Suppose that $D_{i}=1$. (E.11) implies $s_{1i}\geq t_{i}$. Then, (E.9) implies $s_{1i}=t_{1}$, and thus we have $s_{1i}/t_{1}=1=D_{i}$. Next, suppose that $D_{i}=0$. (E.10) and (E.12) imply $s_{1i}=0$. By (E.13), we have $s_{1i}/t_{1}=0=D_{i}$.
The problem corresponding to (E.6) can be formulated in a similar manner.
Appendix F More detail about the legal framework
F.1. Title VII of the Civil Rights Act of 1964
The exact legal process for evaluating a disparate impact case depends on the context and in particular the law prohibiting the behavior. Different laws provide different rights, remedies, and standards of evidence. However, there is a general framework that is commonly used as a blueprint. The framework was developed in the context of Title VII of the Civil Rights Act of 1964, which concerns employment discrimination. [^19] Historically, the framework was introduced by the US Supreme Court in the case Griggs v. Duke Power Co. (1971). The Civil Rights Act of 1964 was later amended by the Civil Rights Act of 1991 to codify the framework. In particular, the law prohibits employers from using “a facially neutral employment practice that has an unjustified adverse impact on members of a protected class. A facially neutral employment practice is one that does not appear to be discriminatory on its face; rather it is one that is discriminatory in its application or effect.”
The process itself is outlined in 42 U.S. Code § 2000e–2(k) which reads
  • (k)
    Burden of proof in disparate impact cases
    • (1)
      (A) An unlawful employment practice based on disparate impact is established under this subchapter only if—
      • (i)
        a complaining party demonstrates that a respondent uses a particular employment practice that causes a disparate impact on the basis of race, color, religion, sex, or national origin and the respondent fails to demonstrate that the challenged practice is job related for the position in question and consistent with business necessity; or
      • (ii)
        the complaining party makes the demonstration described in subparagraph (C) with respect to an alternative employment practice and the respondent refuses to adopt such alternative employment practice.
      • (i)
        With respect to demonstrating that a particular employment practice causes a disparate impact as described in subparagraph (A)(i), the complaining party shall demonstrate that each particular challenged employment practice causes a disparate impact, except that if the complaining party can demonstrate to the court that the elements of a respondent’s decision making process are not capable of separation for analysis, the decision making process may be analyzed as one employment practice.
      • (ii)
        If the respondent demonstrates that a specific employment practice does not cause the disparate impact, the respondent shall not be required to demonstrate that such practice is required by business necessity.
      (C) The demonstration referred to by subparagraph (A)(ii) shall be in accordance with the law as it existed on June 4, 1989, with respect to the concept of “alternative employment practice”.
    • (2)
      A demonstration that an employment practice is required by business necessity may not be used as a defense against a claim of intentional discrimination under this subchapter.
    • (3)
      Notwithstanding any other provision of this subchapter, a rule barring the employment of an individual who currently and knowingly uses or possesses a controlled substance, as defined in schedules I and II of section 102(6) of the Controlled Substances Act (21 U.S.C. 802(6)), other than the use or possession of a drug taken under the supervision of a licensed health care professional, or any other use or possession authorized by the Controlled Substances Act [21 U.S.C. 801 et seq.] or any other provision of Federal law, shall be considered an unlawful employment practice under this subchapter only if such rule is adopted or applied with an intent to discriminate because of race, color, religion, sex, or national origin.
This process was originally designed to be used by parties seeking judicial enforcement, and so is adversarial. For example, the complaining party may be an employee who is asking a judge to prohibit a potentially discriminatory employment practice and the respondent is the employer implementing the practice. In this case, the judge would evaluate the three parts of the process (the prima facie case, the business necessity defense, and the existence of suitable alternatives), and then determine whether the practice is prohibited under Title VII.
F.2. Title VI of the Civil Rights Act of 1964
Many policies of interest to economists are implemented by organizations that receive federal funding. For example, many hospitals receive funding from Medicare or Hill-Burton. Because these organizations receive federal funding, their policies are prohibited from having a disparate impact by Title VI of the Civil Rights Act of 1964. This prohibition is described in 28 C.F.R. § 42.104(b)(2) which reads
  • (2)
    A recipient, in determining the type of disposition, services, financial aid, benefits, or facilities which will be provided under any such program, or the class of individuals to whom, or the situations in which, such will be provided under any such program, or the class of individuals to be afforded an opportunity to participate in any such program, may not, directly or through contractual or other arrangements, utilize criteria or methods of administration which have the effect of subjecting individuals to discrimination because of their race, color, or national origin, or have the effect of defeating or substantially impairing accomplishment of the objectives of the program as respects individuals of a particular race, color, or national origin.
A key difference between that of a Title VII disparate impact case, is that in the setting of Title VI there is no private right to action (see, specifically, Alexander v. Sandoval, 2001). That is, individuals cannot file civil actions relying on the Title VI disparate impact standard. Instead, it is the responsibility of the funding agency to enforce the law, which can take place both by investigating individual complaints and conducting compliance reviews. However, to fulfill this responsibility, the Department of Justice still recommends using the process outlined for Title VII described above. In particular, the Department of Justice Civil Rights Division’s Title VI Manual (available at https://www.justice.gov/crt/fcs/T6manual) writes:
“In administrative investigations, this court-developed burden shifting framework serves as a useful paradigm for organizing the evidence. Agency investigations, however, often follow a non-adversarial model in which the agency collects all relevant evidence then determines whether the evidence establishes discrimination. Under this model, agencies often do not shift the burdens between complainant and recipient when making findings. For agencies using this method, the following sections serve as a resource for conducting an investigation and developing an administrative enforcement action where appropriate.”
In addition, while the Title VI Manual identifies the funding agency as ultimately responsible for evaluating less discriminatory alternatives, it also writes that “Although agencies bear the burden of evaluating less discriminatory alternatives, agencies sometimes impose additional requirements on recipients to consider alternatives before taking action. These requirements can affect the legal framework by requiring recipients to develop the evidentiary record related to alternatives as a matter of course, before and regardless of whether an administrative complaint is even filed. Such requirements recognize that the recipient is in the best position to complete this task, having the best understanding of its goals, and far more ready access to the information necessary to identify alternatives and conduct a meaningful analysis.”
Appendix G Supplementary Material to Section 5
G.1. Robustness checks related to selection rule choice
We report here the point estimates of accuracy and fairness for both the candidate and status quo algorithms, as well as the $p$-values that emerge from our statistical test of the null hypothesis that the status quo algorithm is not strictly FA-dominated, for LASSO (Table 2) and random forest (Table 3). The results are nearly identical to those reported in Table 1 for linear regression.
Accuracy (Black) Accuracy (White) Unfairness
$a_{1}$ $a_{0}$ $p_{b}$ $a_{1}$ $a_{0}$ $p_{w}$ $a_{1}$ $a_{0}$ $p_{f}$ $p$
Iteration 1 7.17 6.33 0.0004 7.25 5.15 0.0000 0.09 1.17 0.0000 0.0004
Iteration 2 7.38 6.37 0.0001 7.25 5.14 0.0000 0.13 1.23 0.0000 0.0001
Iteration 3 7.64 6.67 0.0000 7.31 5.17 0.0000 0.34 1.50 0.0000 0.0000
Iteration 4 7.21 6.38 0.0002 7.28 5.13 0.0000 0.07 1.25 0.0000 0.0002
Iteration 5 7.63 6.88 0.0022 7.36 5.28 0.0000 0.27 1.60 0.0000 0.0022
Table 2. The candidate algorithm $a_{1}$ in the table is based on LASSO. Reported $p$-values are computed via bootstrap with 10,000 iterations. The median $p$-value $p_{\mathrm{med}}$ is 0.0002.
Accuracy (Black) Accuracy (White) Unfairness
$a_{1}$ $a_{0}$ $p_{b}$ $a_{1}$ $a_{0}$ $p_{w}$ $a_{1}$ $a_{0}$ $p_{f}$ $p$
Iteration 1 7.17 6.33 0.0006 7.18 5.15 0.0000 0.01 1.17 0.0000 0.0006
Iteration 2 7.39 6.37 0.0001 7.21 5.14 0.0000 0.19 1.23 0.0001 0.0001
Iteration 3 7.55 6.67 0.0000 7.25 5.17 0.0000 0.30 1.50 0.0000 0.0000
Iteration 4 7.30 6.38 0.0000 7.08 5.13 0.0000 0.22 1.25 0.0001 0.0001
Iteration 5 7.66 6.88 0.0017 7.28 5.28 0.0000 0.38 1.60 0.0000 0.0017
Table 3. The candidate algorithm $a_{1}$ in the table is based on random forest. Reported $p$-values are computed via bootstrap with 10,000 iterations. The median $p$-value $p_{\mathrm{med}}$ is 0.0001.
G.2. Robustness checks related to program effect
In our baseline analysis, we use the total number of active chronic illnesses in the year following enrollment in the management program as the outcome variable $Y$. As a robustness check, here we repeat this analysis using the total number of active chronic illnesses in the same year as enrollment in the management program. We report point estimates of accuracy and fairness for both the candidate and status quo algorithms, along with corresponding $p$-values. The results are consistent with those observed in our baseline analysis: we reject the null hypothesis that it is not possible to reduce the algorithm’s disparate impact without compromising on the accuracy of its predictions. For additional robustness checks related to the size of the program effect, see the supplementary materials of Obermeyer et al. (2019).
Figure 6. We report the average value of $U^{g}$ (i.e., number of active chronic diseases in the same year as the enrollment decision, conditional on automatic enrollment) for each group $g$ across the $K=5$ iterations of our procedure. This figure is qualitatively similar to Figure 2.
Accuracy (Black) Accuracy (White) Unfairness
$a_{1}$ $a_{0}$ $p_{b}$ $a_{1}$ $a_{0}$ $p_{w}$ $a_{1}$ $a_{0}$ $p_{f}$ $p$
Iteration 1 9.39 7.43 0.0000 8.97 5.82 0.0000 0.41 1.60 0.0000 0.0000
Iteration 2 9.18 7.49 0.0001 8.86 5.85 0.0000 0.32 1.64 0.0000 0.0001
Iteration 3 9.61 7.82 0.0000 9.08 5.95 0.0000 0.53 1.87 0.0000 0.0000
Iteration 4 9.33 7.46 0.0000 8.81 5.76 0.0000 0.52 1.70 0.0000 0.0000
Iteration 5 9.64 8.14 0.0012 8.98 5.94 0.0000 0.67 2.20 0.0000 0.0012
Table 4. The candidate algorithm $a_{1}$ in the table is based on linear regression. Reported $p$-values are computed via bootstrap with 10,000 iterations. The median $p$-value $p_{\mathrm{med}}$ is smaller than 0.0001.
Accuracy (Black) Accuracy (White) Unfairness
$a_{1}$ $a_{0}$ $p_{b}$ $a_{1}$ $a_{0}$ $p_{w}$ $a_{1}$ $a_{0}$ $p_{f}$ $p$
Iteration 1 9.26 7.43 0.0004 9.01 5.82 0.0000 0.24 1.60 0.0000 0.0004
Iteration 2 9.17 7.49 0.0001 8.86 5.85 0.0000 0.30 1.64 0.0000 0.0001
Iteration 3 9.46 7.82 0.0000 9.11 5.95 0.0000 0.35 1.87 0.0000 0.0000
Iteration 4 9.27 7.46 0.0002 8.83 5.76 0.0000 0.44 1.70 0.0000 0.0002
Iteration 5 9.51 8.14 0.0022 9.01 5.94 0.0000 0.50 2.20 0.0000 0.0022
Table 5. The candidate algorithm $a_{1}$ in the table is based on LASSO. Reported $p$-values are computed via bootstrap with size 10,000. The median $p$-value $p_{\mathrm{med}}$ is 0.0002.
Accuracy (Black) Accuracy (White) Unfairness
$a_{1}$ $a_{0}$ $p_{b}$ $a_{1}$ $a_{0}$ $p_{w}$ $a_{1}$ $a_{0}$ $p_{f}$ $p$
Iteration 1 8.97 7.43 0.0006 8.47 5.82 0.0000 0.51 1.60 0.0000 0.0006
Iteration 2 8.90 7.49 0.0001 8.33 5.85 0.0000 0.57 1.64 0.0001 0.0001
Iteration 3 8.93 7.82 0.0000 8.65 5.95 0.0000 0.28 1.87 0.0000 0.0000
Iteration 4 8.78 7.46 0.0000 8.40 5.76 0.0000 0.37 1.70 0.0001 0.0001
Iteration 5 9.17 8.14 0.0017 8.48 5.94 0.0000 0.69 2.20 0.0000 0.0017
Table 6. The candidate algorithm $a_{1}$ in the table is based on random forest. Reported $p$-values are computed via bootstrap with size 10,000. The median $p$-value $p_{\mathrm{med}}$ is 0.0001.